### v2

# Condensed Interior-Point Methods for Scalable Nonlinear Programming on GPUs

Abstract
This paper explores two variants of condensed-space interior-point methods designed for GPUs—HyKKT and LiftedKKT—by analyzing their numerical properties through error analysis and assessing their real-world performance via extensive numerical experiments with a fully GPU-resident software implementation. Traditional implementations of interior-point methods (IPMs) involve solving indefinite augmented KKT systems repeatedly by utilizing direct sparse solvers based on the LBL factorization with sophisticated numerical pivoting strategies. While this method achieves high performance and robustness on CPUs, the serial nature of numerical pivoting presents challenges for effective implementation on GPUs. Recently, multiple condensed-space IPM strategies have emerged to address this issue by transforming the KKT system into a symmetric positive-definite matrix, which is more suitable for factorization on GPUs. However, these methods have been demonstrated only on optimal power flow instances, and their numerical properties are currently inadequately understood. In this study, we demonstrate that although the condensed systems show increased ill-conditioning, the inherent structures of the condensed KKT system effectively counterbalance potential accuracy loss in the IPM. Furthermore, we provide numerical results that thoroughly assess the capabilities of a fully GPU-resident nonlinear programming software stack, comprising MadNLP.jl (a filter line-search IPM solver), cuDSS (a direct sparse solver leveraging Cholesky factorization), and ExaModels.jl (a modeling framework), by benchmarking their performance against the pglib-opf and CUTEst libraries. Our findings suggest that the GPU framework holds promise for solving highly sparse large-scale nonlinear programs, such as optimal power flow instances, except for diminished robustness and limited speedups for edge cases observed within CUTEst instances.
1 Introduction
General-purpose GPU computing has emerged as a fundamental component of scientific computing and machine learning, particularly in its ability to train and deploy large-scale artificial intelligence (AI) models. As a computing platform, the modern GPU hardware offer two major advantages: (1) substantial parallel computing power and (2) improved power efficiency.
Despite their achievements in scientific computing and machine learning, GPUs have been underutilized in classical mathematical programming, including linear programming (LP) and nonlinear programming (NLP). In machine learning applications, the parallelization is mostly utilized at the low-level matrix operations, via the libraries like NVIDIA’s cuBLAS or AMD’s rocBLAS. In contrast, most classical mathematical programming applications exhibit limited parallelism at the BLAS operation level due to the inherent sparsity of the underlying problem instances, which prevents straightforward utilization of parallel resources. Although certain categories of sparse matrix operations can be parallelized, the development of efficient sparse matrix kernels on GPUs poses significant challenges, particularly for sparse matrix factorizations. Classical solution methodologies typically depend on direct linear solvers that employ sophisticated numerical pivoting strategies to maintain numerical stability. However, these techniques are inherently sequential, creating difficulties for effective parallelization on GPUs. As a result, until recently, the full potential of GPU computing has not been realized within the mathematical programming domain.
Recent advancements in GPU computing techniques are beginning to change this landscape. Here are some significant developments, driven either by the hardware, software, or the emergence of new applications:
  1. 1.
    Development of mature direct sparse linear solvers: Recent developments in GPU-based sparse direct solvers, such as NVIDIA’s cuDSS nvidiaNVIDIACuDSSPreview , have significantly improved the performance of the direct, repetitive solution of sparse linear systems. These solvers now offer performance that rivals or exceeds that of traditional CPU-based solvers, though the types of factorizations currently supported are limited to Cholesky, $\mathrm{LDL^{T}}$, and LU as of now nvidiaNVIDIACuDSSPreview ; shin2023accelerating .
  2. 2.
    Progress in automatic differentiation: GPUs are well-suited for automatic differentiation, a key component of both scientific simulations and machine learning jax2018github ; enzyme2021 . Efficient implementations for algebraic modeling and derivative evaluation on GPUs have now become available. For large-scale problem instances with highly repetitive patterns, the derivative evaluation can be performed at speeds that are orders of magnitude faster than their CPU counterparts shin2023accelerating . Notably, derivative evaluation kernels can be created in a fully automated fashion. This removes the barrier of having to write custom derivative evaluation kernels for each problem instance, which has been a major bottleneck in the past.
  3. 3.
    Rise of GPU-dominated HPC: With the emergence of exascale supercomputers (e.g., Frontier, Aurora), GPU support has become a core requirement for utilizing these HPC resources in mathematical optimization applications.
  4. 4.
    Growing interest in batched optimization: In machine learning, embedding optimization problems into neural networks requires solving many small problems in parallel. This has led to the development of GPU-native batched solvers capable of handling thousands of small problems simultaneously amos2017optnet ; pineda2022theseus . Similar ideas have been applied to decomposition methods in large-scale optimization kimLeveragingGPUBatching2021 , but extending these methods to large-scale problems remains difficult, as current implementations often rely on dense linear algebra.
Due to this renewed interest in GPU-based optimization, there has been a surge of research in this area, particularly in the context of developing scalable algorithms and software implementations that can harness the power of GPUs. Here, we provide a general overview of the state of the art in GPU-based optimization, highlighting the challenges and opportunities in this area. The recent developments in GPU-based optimization can be broadly categorized into three main approaches: (i) first-order methods, (ii) second-order methods with iterative linear algebra, and (iii) second-order methods with sparse direct solvers. In terms of numerical implementation, the first two approaches rely mostly on repetitive sparse matrix-vector multiplications, while the third approach requires sparse direct solvers.
First-order methods for large-scale optimization.
Although first-order methods have been extensively used for large-scale optimization, especially in machine learning applications, their application to classical mathematical programming remains limited due to their slow linear convergence rate. However, for extremely large-scale instances of linear programs, where the numerical factorization of the KKT systems or normal equations is infeasible, first-order methods can be a promising alternative. Recently, a special type of primal-dual first-order method, known as the primal-dual hybrid gradient (PDHG) method, has been developed to target such extremely large problems. This method is particularly effective for large-scale linear programs, as it can handle problems with billions of variables and constraints. This algorithm is implemented within a solver called PDLP, and it is adapted to GPU architectures lu2023cupdlp ; lu2023cupdlp2 . Recently, this algorithm has been enhanced with the restarted Halpern PDHG method with reflection and a PID-controlled primal weight update strategy, and has been shown to achieve greater speed luCuPDLPFurtherEnhanced2025 . Although these methods have faced challenges in obtaining high precision solutions, for extremely large-scale instances of linear programs, they have shown to outperform even commercial solvers like Gurobi when fully GPU-accelerated lu2023cupdlp ; lu2023cupdlp2 .
Second-order methods with iterative linear algebra.
An alternative to first-order methods is to perform second-order methods with iterative linear algebra, such as Krylov subspace methods. Krylov subspace methods require only matrix-vector products and are inherently parallelizable. These methods are well-suited for large-scale optimization problems, as they can efficiently solve the KKT systems arising in interior-point methods without requiring full matrix factorizations. However, most second-order algorithms, such as the primal-dual interior-point method, rely on barrier functions to handle inequality constraints, which can lead to ill-conditioned KKT systems; thus, the iterative linear solvers need to be employed with careful KKT system reformulation and/or preconditioning strategies curtisNoteImplementationInteriorpoint2012 ; rodriguezScalablePreconditioningBlockstructured2020 . Promising results have been reported in convex settings schubigerGPUAccelerationADMM2020 by combining the operator splitting methods and preconditioned conjugate gradient. For nonlinear programming, a strategy based on the Augmented Lagrangian method, interior-point method, and preconditioned conjugate gradient method has been proposed cao2016augmented .
Second-order methods for large-scale optimization.
The previous two strategies are well suited to large-scale convex optimization, but extending their success to general nonlinear programming remains challenging. For nonconvex settings, first-order or iterative linear algebra-based implementations often struggle to achieve high accuracy. For instance, it has been previously reported that ADMM fails to converge reliably below a tolerance of $10^{-3}$ on the AC optimal power flow (OPF) problem kimLeveragingGPUBatching2021 . However, second-order methods with direct linear algebra, when implemented correctly, can offer significantly faster convergence and more robust convergence behaviors. By solving a Newton system at each iteration, they can provide fast local convergence and high robustness, even in the presence of nonconvexities.
Second-order methods with direct linear algebra are substantially more challenging to implement on GPUs. This is because direct solution methods require the efficient factorization of large sparse, indefinite Karush–Kuhn–Tucker (KKT) systems. Even as of now, to the best of our knowledge, no GPU-based direct solver can directly handle these sparse indefinite KKT systems encountered in large-scale nonlinear optimization applications. Historically, GPU-based sparse solvers have lagged behind their CPU counterparts in both performance and robustness. Notably, it has been reported that GPU solvers do not offer significant acceleration for power system optimization problems swirydowicz2021linear ; tasseff2019exploring . Although several direct sparse solvers, such as CUSOLVERRF, were available on GPUs, they were not widely adopted in practice for solving large-scale nonlinear optimization problems due to their limited performance and robustness.
This situation has changed in recent years, with two advances: the development of pivoting-free interior-point methods and the emergence of new GPU-native sparse direct solvers. First, the KKT system reformulation strategy, known as condensed-space methods, has been proposed to address the challenges of solving indefinite KKT systems within interior-point methods. There are various forms of such reformulations, including the equality relaxation method (so-called LiftedKKT) and the hybrid Golub-Greif method (so-called HyKKT; adapted in regev2023hykkt ). Alternatively, null-space methods (also known as reduced-Hessian methods) eliminate equality constraints to reduce the KKT system to a dense problem, which maps well to GPU architectures pacaud2022condensed . These strategies transform the indefinite KKT system into a symmetric positive-definite (SPD) system, which can be solved using Cholesky factorization with static pivoting.
Furthermore, recent advances in GPU-based sparse direct solvers have significantly improved their performance and robustness. Especially for solving positive definite or quasi-definite systems, which do not require sophisticated pivoting strategies, GPU solvers now often surpass the performance of CPU-based solvers. Notably, in November 2023, NVIDIA released cuDSS, a new sparse direct solver providing efficient sparse factorizations with significantly improved performance over earlier GPU-based solvers. These advances, combined with recent developments in automatic differentiation and interior-point methods that avoid numerical pivoting, enable robust and scalable second-order methods to be implemented fully on GPUs. Coupling cuDSS with a GPU-native AD engine allows us to solve large-scale nonlinear problems like OPF up to 10× faster than previous state-of-the-art methods shin2023accelerating .
Nevertheless, the condensed-space methods have several known limitations. Although condensed KKT systems remain sparse in many problem instances, performing the elimination of the inequality constraints often leads to significantly increased nonzero entries in the condensed KKT system for certain instances. The worst case is a fully dense constraint where both the equality relaxation and HyKKT methods mentioned above can lead to fully dense condensed KKT systems, leading to a prohibitively expensive to solve. Due to this aspect, condensed-space methods are inherently unsuitable for problems with dense constraints. Accordingly, when these methods are compared against standard benchmark sets, such as CUTEst, they are bound to exhibit varied performance, due to the presence of the instances with such dense constraints. Furthermore, due to the nature of the condensed KKT systems, the condition number of the condensed KKT system can be significantly larger than that of the original KKT system, which can lead to numerical issues.
A full Software Stack for GPU-Accelerated Nonlinear Programming.
The state-of-the-art nonlinear optimization software stack consists of three key components: (i) a high-level modeling interface, (ii) a nonlinear optimization solver, (iii) a sparse linear solver. In the classical CPU-based setting, the most widely used software stack for nonlinear optimization consists of AMPL fourer1990ampl for modeling, Ipopt for the solver, and the HSL library (MA27, MA57, etc.) for sparse factorization. It is important to note that in order to fully benefit from the computational power of GPUs, the entire optimization pipeline must be implemented on the GPU. If any component is not GPU-compatible, it is necessary to introduce costly CPU–GPU communication, similar to double-buffering in MPI-based distributed systems. This avoids expensive transfers between host (CPU) and device (GPU) memory. For older generations of the PCIe interface, this can be a major bottleneck. Thus, in order to achieve high performance, it is essential that all components of the optimization pipeline are GPU-native and operate directly in GPU memory. The GPU-native software stack used within this paper is provided by MadNLP.jl, ExaModels.jl, and cuDSS. This paper focuses on the second component, the nonlinear optimization solver, while the other two components are briefly described below.
ExaModels.jl is an algebraic modeling system designed for efficient model function and derivative evaluations on GPUs shin2023accelerating . ExaModels.jl provides a symbolic modeling front-end where users can define their optimization problems in a structure-revealing fashion, similarly to the constraint template implemented in Gravity hijazi2018gravity . Unlike classical modeling interfaces such as AMPL, JuMP, and CasADi, which provide flexible modeling features, ExaModels.jl follows a different philosophy: it enforces strict compliance with the iterator-based modeling syntax, allowing the modeling system to store the model equations in a structured fashion. This ensures that the model function and derivative evaluations can be implemented as parallelizable GPU kernels. For the structured model functions, ExaModels.jl applies the reverse-mode automatic differentiation to compute the derivatives, then executed over GPU-resident data arrays. To the best of our knowledge, ExaModels.jl is currently the only available GPU-native sparse modeling interface that provides automatic differentiation capabilities for nonlinear programming beyond unconstrained problems in machine learning.
NVIDIA’s cuDSS library nvidiaNVIDIACuDSSPreview provides a high-performance sparse direct solver. The library supports Cholesky, LDL, and LU factorizations, enabling the efficient solution of symmetric positive definite and quasi-definite systems. When the hybrid host/device execution mode is disabled, reordering (a major part of the analysis phase) is executed on the host, while symbolic factorization (another part of the analysis phase), numerical factorization, and solve are executed on the GPU. Since ordering is inherently sequential, it is precomputed once on the CPU and transferred to the GPU without any impact on performance. Partial pivoting strategies are available, but the 2x2 pivoting strategy — commonly employed within sparse indefinite solvers – is not supported, thus limiting their direct usage within the classical augmented system-based interior-point methods. CUDSS.jl, a community-maintained Julia wrapper for cuDSS, provides a user-friendly interface to the library, enabling seamless integration with Julia-based optimization solvers such as MadNLP.jl.
1.1 Contributions
In this article, we aim to theoretically and empirically assess the capabilities of the GPU-accelerated solution strategies for large-scale sparse nonlinear programs. Although we utilize the full GPU-resident software stack, we focus on the nonlinear optimization solver component, which is based on filter line-search interior point method with two variants of condensed-space KKT system reformulation strategies: HyKKT regev2023hykkt and LiftedKKT shin2023accelerating . Our main contributions are summarized below:
  1. 1.
    We perform a rigorous error analysis of the condensed-space IPM methods for two condensed-space methods under study: HyKKT and LiftedKKT. We extend classical results from wright1998ill to show that the condensed KKT matrices are structurally ill-conditioned in a controlled manner. This yields accurate error bounds for the Schur complement used in the HyKKT method as well as for the condensed KKT system within LiftedKKT.
  2. 2.
    We provide a fully GPU-resident software stack—consisting of ExaModels.jl (derivative evaluation), MadNLP.jl (interior point solver), and NVIDIA cuDSS —for solving large-scale sparse nonlinear optimization problems, based on two different condensed-space KKT system reformulation strategies: HyKKT and LiftedKKT. Compared to our previous work shin2023accelerating , we (i) have additionally implemented the HyKKT method, and (ii) have replaced the linear solver cusolverRF used in shin2023accelerating with NVIDIA’s new cuDSS library. As a result, the interior-point method is executed entirely on the GPU, from model evaluation to KKT system solving.
  3. 3.
    We conduct a comparative study of the two condensed-space methods, analyzing their accuracy and runtime performance. We go beyond traditional OPF test cases by including large-scale problems from the CUTEst benchmark gould2015cutest . Our findings demonstrate that the condensed-space IPM enables a remarkable tenfold acceleration in solving large-scale OPF instances when utilizing the GPU. However, performance outcomes on the CUTEst benchmark exhibit more variability. For larger problems (where Ipopt takes longer than 1 second), our GPU solver is competitive with the CPU-based Ipopt solver, but exhibits less robustness and performance variability across different problem types. One of the main reasons for variability in performance is due to the fact that a single dense row can incur a dense condensed KKT system, which can make the factorization of the condensed KKT system significantly more expensive.
  4. 4.
    We demonstrate that condensed-space methods can be beneficial for a wide range of parallel architectures, not limited to GPUs. In fact, the condensed-space strategies presented in this paper are suitable not only for GPUs but also for other parallel architectures, such as multi-core CPUs on shared memory. We provide numerical results of the two condensed-space methods on multi-core CPUs, showing that they can offer substantial speed-ups for large-scale instances, though less pronounced than on GPUs. In particular, our numerical results reveal that the parallel CPU solvers, such as HSL MA86 and Pardiso, are particularly efficient at solving the condensed KKT systems. On the largest OPF instance, we observe that the parallel CPU solvers are almost 5 times faster than HSL MA27, whereas NVIDIA cuDSS is itself 25 times faster than HSL MA27.
1.2 Notations
By default, the norm $\|\cdot\|$ refers to the 2-norm. We define the conditioning of a matrix $A$ as $\kappa_{2}(A)=\|A\|\|A^{-1}\|$. For any real number $x$, we denote by $\widehat{x}$ its floating point representation. We denote $\mathbf{u}$ as the smallest positive number such that $\widehat{x}\leq(1+\tau)x$ for $|\tau|<\mathbf{u}$. In double precision, $\mathbf{u}=1.1\times 10^{-16}$. We use the following notations to proceed with our error analysis. For $p\in\mathbb{N}$ and a positive variable $h$:
  • We write $x=O(h^{p})$ if there exists a constant $b>0$ such that $\|x\|\leq bh^{p}$;
  • We write $x=\Omega(h^{p})$ if there exists a constant $a>0$ such that $\|x\|\geq ah^{p}$;
  • We write $x=\Theta(h^{p})$ if there exist two constants $0<a<b$ such that $ah^{p}\leq\|x\|\leq bh^{p}$.
2 Primal-dual interior-point method
The interior-point method (IPM) is among the most popular algorithms to solve nonlinear programs. The basis of the algorithm is to reformulate the Karush-Kuhn-Tucker (KKT) conditions of the nonlinear program as a smooth system of nonlinear equations using a homotopy method nocedal_numerical_2006 . In a standard implementation, the resulting system is solved iteratively with a Newton method (used in conjunction with a line-search method for globalization).
This section provides a brief overview of the IPM algorithm and their numerical implementation. In Section 2.1, we give a description of a nonlinear programming problem formulation. We detail in Section 2.2 the Newton step computation within each IPM iteration.
2.1 Problem formulation and KKT conditions
We are interested in solving the following nonlinear program:
$\min_{x\in\mathbb{R}^{n}}\;f(x)\quad\text{subject to}\quad g(x)=0\;,~h(x)\leq 0\;,$ (1)
with $f:\mathbb{R}^{n}\to\mathbb{R}$ a real-valued function encoding the objective, $g:\mathbb{R}^{n}\to\mathbb{R}^{m_{e}}$ encoding the equality constraints, and $h:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ encoding the inequality constraints. In what follows, we suppose that the functions $f,g,h$ are smooth and twice differentiable.
We reformulate (1) using non-negative slack variables $s\geq 0$ into the equivalent formulation
$\min_{x\in\mathbb{R}^{n},~s\in\mathbb{R}^{m_{i}}}\;f(x)\quad\text{subject to}\quad\left\{\begin{aligned} &g(x)=0\;,~h(x)+s=0\;,\\ &s\geq 0\;.\end{aligned}\right$. (2)
In (2), the inequality constraints are encoded inside the variable bounds on the slack variables. We denote by $y\in\mathbb{R}^{m_{e}}$ and $z\in\mathbb{R}^{m_{i}}$ the multipliers associated resp. to the equality constraints and the inequality constraints. Similarly, we denote by $v\in\mathbb{R}^{m_{i}}$ the multipliers associated to the bounds $s\geq 0$.
Using the dual variable $(y,z,v)$, we define the Lagrangian of (2) as
$L(x,s,y,z,v)=f(x)+y^{\top}g(x)+z^{\top}\big{(}h(x)+s\big{)}-v^{\top}s\;.$ (3)
The KKT conditions of (2) are:
$\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z=0\;,$ (4a)
$z-v=0\;,$ (4b)
$g(x)=0\;,$ (4c)
$h(x)+s=0\;,$ (4d)
$0\leq s\perp v\geq 0\;.$ (4e)
The notation $s\perp v$ is a shorthand for the complementarity condition $s_{i}v_{i}=0$ (for all $i=1,\cdots,n$). The set of active constraints at a point $x$ is denoted by:
$\mathcal{B}(x):=\{i\in\{1,\cdots,m_{i}\}\;|\;h_{i}(x)=0\}\;.$ (5)
The inactive set is defined as the complement $\mathcal{N}(x):=\{1,\cdots,m_{i}\}\setminus\mathcal{B}(x)$. We note $m_{a}$ the number of active constraints. The active Jacobian is defined as $A(x):=\begin{bmatrix}\nabla g(x)\\ \nabla h_{\mathcal{B}}(x)\end{bmatrix}\in\mathbb{R}^{(m_{e}+m_{a})\times n}$.
2.2 Solving the KKT conditions with the interior-point method
The interior-point method aims at finding a stationary point satisfying the KKT conditions (4). The complementarity constraints (4e) render the KKT conditions non-smooth, complicating the solution of the whole system (4). IPM uses a homotopy continuation method to solve a simplified version of (4), parameterized by a barrier parameter $\mu>0$ (nocedal_numerical_2006, , Chapter 19). For positive $(x,s,v)>0$, we solve the system
$F_{\mu}(x,s,y,z,v)=\begin{bmatrix}\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z\\ z-v\\ g(x)\\ h(x)+s\\ Sv-\mu e\end{bmatrix}=0\;.$ (6)
We introduce in (6) the diagonal matrices $X=\operatorname{diag}(x_{1},\cdots,x_{n})$ and $S=\operatorname{diag}(s_{1},\cdots,s_{m_{i}})$, along with the vector of ones $e$. As we drive the barrier parameter $\mu$ to 0, the solution of the system $F_{\mu}(x,s,y,z,v)=0$ tends to the solution of the KKT conditions (4).
We note that at a fixed parameter $\mu$, the function $F_{\mu}(\cdot)$ is smooth. Hence, the system (6) can be solved iteratively using a regular Newton method. For a primal-dual iterate $w_{k}:=(x_{k},s_{k},y_{k},z_{k},v_{k})$, the next iterate is computed as $w_{k+1}=w_{k}+\alpha_{k}d_{k}$, where $d_{k}$ is a descent direction computed by solving the linear system
$\nabla_{w}F_{\mu}(w_{k})d_{k}=-F_{\mu}(w_{k})\;.$ (7)
The step $\alpha_{k}$ is computed using a line-search algorithm, in a way that ensures that the bounded variables remain positive at the next primal-dual iterate: $(x_{k+1},s_{k+1},v_{k+1})>0$. Once the iterates are sufficiently close to the central path, the IPM decreases the barrier parameter $\mu$ to find a solution closer to the original KKT conditions (4).
In IPM, the bulk of the workload is the computation of the Newton step (7), which involves assembling the Jacobian $\nabla_{w}F_{\mu}(w_{k})$ of the KKT conditions and solving the resulting linear system to compute the descent direction $d_{k}$. By writing out all the blocks, the system in (7) expands as the $5\times 5$ unreduced KKT system of the following form:
$\begin{bmatrix}W_{k}&0&G_{k}^{\top}&H_{k}^{\top}&\phantom{-}0\\ 0&0&0\phantom{{}^{\top}}&I\phantom{{}^{\top}}&-I\\ G_{k}&0&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0\\ H_{k}&I&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0\\ 0&V_{k}&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}S_{k}\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{s}\\ d_{y}\\ d_{z}\\ d_{v}\end{bmatrix}=-\begin{bmatrix}\nabla_{x}L(w_{k})\\ z_{k}-v_{k}\\ g(x_{k})\\ h(x_{k})+s_{k}\\ S_{k}v_{k}-\mu e\end{bmatrix}\;$, ($K_{3}$)
where we have introduced the Hessian $W_{k}=\nabla^{2}_{xx}L(w_{k})$ and the two Jacobians $G_{k}=\nabla g(x_{k})$, $H_{k}=\nabla h(x_{k})$. In addition, we define $X_{k}$, $S_{k}$, $U_{k}$ and $V_{k}$ the diagonal matrices built respectively from the vectors $x_{k}$, $s_{k}$, $u_{k}$ and $v_{k}$. Note that ($K_{3}$) can be symmetrized by performing simple block row and column operations. In what follows, we will omit the iteration counter $k$ to simplify the notations.
Augmented KKT system.
It is usual to remove in ($K_{3}$) the blocks associated to the bound multipliers $v$ and solve instead the regularized $4\times 4$ symmetric system, called the augmented KKT system:
$\begin{bmatrix}W+\delta_{w}I&0&\phantom{-}G^{\top}&\phantom{-}H^{\top}\\ 0&D_{s}+\delta_{w}I&\phantom{-}0\phantom{{}^{\top}}&\phantom{-}I\phantom{{}^{\top}}\\ G&0&-\delta_{c}I&\phantom{-}0\phantom{{}^{\top}}\\ H&I&\phantom{-}0\phantom{{}^{\top}}&-\delta_{c}I\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{s}\\ d_{y}\\ d_{z}\end{bmatrix}=-\begin{bmatrix}r_{1}\\ r_{2}\\ r_{3}\\ r_{4}\end{bmatrix}\;$, ($K_{2}$)
with the diagonal matrix $D_{s}:=S^{-1}V$. The vectors forming the right-hand-sides are given respectively by $r_{1}:=\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z$, $r_{2}:=z-\mu S^{-1}e$, $r_{3}:=g(x)$, $r_{4}:=h(x)+s$. Once ($K_{2}$) is solved, we recover the updates on bound multipliers with $d_{v}=-S^{-1}(Vd_{s}-\mu e)-v$.
Note that we have added additional regularization terms $\delta_{w}\geq 0$ and $\delta_{c}\geq 0$ in ($K_{2}$), to ensure the matrix is invertible. Without the regularization terms in ($K_{2}$), the augmented KKT system is non-singular if and only if the Jacobian $J=\begin{bmatrix}G\;&\;0\\ H\;&\;I\end{bmatrix}$ is full row-rank and the matrix $\begin{bmatrix}W&0\\ 0&D_{s}\end{bmatrix}$ projected onto the null-space of the Jacobian $J$ is definite benzi2005numerical . The condition is satisfied if the inertia (defined as the respective numbers of positive, negative and zero eigenvalues) of the matrix ($K_{2}$) is $(n+m_{i},m_{i}+m_{e},0)$. We use the inertia-controlling method introduced in (wachter2006implementation, , Section 3.1) to regularize the augmented matrix by adding multiple of the identity on the diagonal of ($K_{2}$) if the inertia is not equal to $(n+m_{i},m_{e}+m_{i},0)$.
As a consequence, the system ($K_{2}$) is usually factorized using an inertia-revealing $\mathrm{LBL^{T}}$ factorization duff1983multifrontal . Krylov methods are often not competitive when solving ($K_{2}$), as the block diagonal terms $D_{s}$ are getting increasingly ill-conditioned near the solution. Their use in IPM has been limited to linear and convex quadratic programming gondzio-2012 (when paired with a suitable preconditioner). We also refer to cao2016augmented for an efficient implementation of a preconditioned conjugate gradient on GPU, for solving the Newton step arising in an augmented Lagrangian interior-point approach.
Condensed KKT system.
The $4\times 4$ KKT system ($K_{2}$) can be further reduced down to a $2\times 2$ system by eliminating the two blocks $(d_{s},d_{z})$ associated to the inequality constraints. The resulting system is called the condensed KKT system:
$\begin{bmatrix}K&\phantom{-}G^{\top}\\ G&-\delta_{c}I\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{y}\end{bmatrix}=-\begin{bmatrix}r_{1}+H^{\top}(D_{H}r_{4}-Cr_{2})\\ r_{3}\end{bmatrix}=:\begin{bmatrix}\bar{r}_{1}\\ \bar{r}_{2}\end{bmatrix}\;$, ($K_{1}$)
where we have introduced the condensed matrix $K:=W+\delta_{w}I+H^{\top}D_{H}H$ and the two diagonal matrices
$C:=\big{(}I+\delta_{c}(D_{s}+\delta_{w}I)\big{)}^{-1}\;,\quad D_{H}:=(D_{s}+\delta_{w}I)C\;.$ (8)
Using the solution of the system ($K_{1}$), we recover the updates on the slacks and inequality multipliers with $d_{z}=-Cr_{2}+D_{H}(Hd_{x}+r_{4})$ and $d_{s}=-(D_{s}+\delta_{w}I)^{-1}(r_{2}+d_{z})$. Using Sylvester’s law of inertia, we can prove that
$\operatorname{inertia}(K_{2})=(n+m_{i},m_{e}+m_{i},0)\iff\operatorname{inertia}(K_{1})=(n,m_{e},0)\;.$ (9)
This indicates that one can check the inertia of $K_{1}$ to indirectly check the inertia of $K_{2}$, which, in turn, can be used to determine the regularization parameters $\delta_{w}$ and $\delta_{c}$ needed to ensure that the solution to the KKT system $d_{k}$ is in a descent direction.
Iterative refinement.
Compared to ($K_{3}$), the diagonal matrix $D_{s}$ introduces an additional ill-conditioning in ($K_{2}$), which is amplified in the condensed form ($K_{1}$): the elements in the diagonal tend to infinity if a variable converges to its bound and to 0 if the variable is inactive. To address the numerical errors arising from such ill-conditioning, most implementations of IPM employ Richardson iterations on the original system ($K_{3}$) to refine the solution returned by the direct sparse linear solver (see (wachter2006implementation, , Section 3.10)).
2.3 Discussion
We have obtained three different formulations of the KKT systems appearing at each IPM iteration. The original formulation ($K_{3}$) has better conditioning than the two alternatives ($K_{2}$) and ($K_{1}$), but it is of much larger size. The second formulation ($K_{2}$), which is often called augmented KKT system, is used by default in state-of-the-art nonlinear solvers wachter2006implementation ; waltz2006interior . The system ($K_{2}$) is typically factorized using a $\mathrm{LBL^{T}}$ factorization: for sparse matrices, the Duff and Reid multifrontal algorithm duff1983multifrontal is the preferred method (as implemented in the HSL linear solvers MA27 and MA57 duff2004ma57 ). The condensed KKT system ($K_{1}$) is often discarded, as its conditioning is worse than that of ($K_{2}$) (implying less accurate solutions). Additionally, condensation may result in increased fill-in within the condensed system ($K_{1}$(nocedal_numerical_2006, , Section 19.3, p.571). In the worst cases, ($K_{1}$) itself may become fully dense if an inequality row is completely dense. Consequently, condensation methods are not commonly utilized in practical optimization settings. To the best of our knowledge, Knitro waltz2006interior is the only solver that supports computing the descent direction with ($K_{1}$).
3 Solving KKT Systems on the GPU
The GPU has emerged as a prominent computing hardware not only for graphics-related tasks but also for general-purpose computing. GPUs employ a SIMD formalism that yields excellent throughput for parallelizing small-scale operations. However, their utility remains limited when computational algorithms require global communication. Sparse factorization algorithms, which heavily rely on numerical pivoting, pose significant challenges for implementation on GPUs. Previous research has demonstrated that GPU-based linear solvers significantly lag behind their CPU counterparts swirydowicz2021linear ; tasseff2019exploring . One emerging strategy is to utilize sparse factorization techniques that do not necessitate numerical pivoting regev2023hykkt ; shin2023accelerating by leveraging the structure of the condensed KKT system ($K_{1}$).
We present two alternative methods to solve ($K_{1}$). On the one hand, HyKKT is introduced in §3.1 and uses the hybrid strategy of Golub & Greif golub2003solving ; regev2023hykkt . On the other hand, LiftedKKT shin2023accelerating employs an equality relaxation strategy and is presented in §3.2. Before we introduce these methods, we make the following observation, which allows us to simplify the setting for our analysis.
Remark 1
The primal-dual regularization parameters $(\delta_{w},\delta_{c})$ are typically determined by examining the inertia of the system $K_{2}$. In particular, the dual regularization $\delta_{c}$ is activated if zero eigenvalues are detected or if the filter line-search algorithm is entering a feasibility restoration phase (wachter2006implementation, , Section 3.3). The primal regularization $\delta_{w}$ is activated if the number of positive eigenvalues is less than the desired amount of $n+m_{i}$. For a positive $\delta_{c}$, one can solve the KKT system ($K_{1}$) by solving the following equations:
$\left(K+\frac{1}{\delta_{c}}G^{\top}G\right)d_{x}=\bar{r}_{1}-\frac{1}{\delta_{c}}\bar{r}_{2}\;,$ (10)
and by setting $d_{y}=\delta_{c}^{-1}(Gd_{x}-\bar{r}_{2})$. The system in (10) is a Schur complement of the original condensed system in ($K_{1}$) and is positive definite if the system in ($K_{1}$) has the correct inertia $(n,m_{e},0)$. That is, under the application of the standard inertia correction procedure, it can always be Cholesky factorizable. Thus, when discussing the special treatment of the condensed KKT system ($K_{1}$), we assume that the dual regularization is $\delta_{c}=0$, as $\delta_{c}>0$ will only simplify the solution procedure.
3.1 Golub & Greif strategy: HyKKT
The Golub & Greif golub2003solving strategy reformulates the KKT system using an Augmented Lagrangian formulation. It has been recently revisited in regev2023hykkt to solve the condensed KKT system ($K_{1}$) on the GPU. The key idea is to reformulate the condensed KKT system ($K_{1}$) in an equivalent form:
$\begin{bmatrix}K_{\gamma}&G^{\top}\\ G&0\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{y}\end{bmatrix}=\begin{bmatrix}\bar{r}_{1}+\gamma G^{\top}\bar{r}_{2}\\ \bar{r}_{2}\end{bmatrix}\;$, (11)
where we introduce the regularized matrix $K_{\gamma}:=K+\gamma G^{\top}G$ (recall $\delta_{c}=0$ is assumed per Remark 1) . Let $\begin{bmatrix}Y&Z\end{bmatrix}$ be an orthogonal matrix, where $Z$ encodes the basis of the null-space of the Jacobian $G$. Using a classical result from debreu1952definite , if $G$ is full row-rank, then there exists a threshold value $\underline{\gamma}$ such that for all $\gamma>\underline{\gamma}$, the reduced Hessian $Z^{\top}KZ$ is positive definite if and only if $K_{\gamma}$ is positive definite. Using Sylvester’s law of inertia stated in (9), we deduce that for $\gamma>\underline{\gamma}$, $\operatorname{inertia}(K_{2})=(n+m_{i},m_{e}+m_{i},0)$ if and only if the condensed matrix $K_{\gamma}$ is positive definite.
The linear solver HyKKT regev2023hykkt leverages the positive definiteness of $K_{\gamma}$ to solve (11) using a hybrid direct-iterative method with the following steps:
  1. 1.
    Assemble $K_{\gamma}$ and factorize it using sparse Cholesky;
  2. 2.
    Solve the Schur complement of (11) using a conjugate gradient (Cg) algorithm to recover the dual descent direction:
    $(GK_{\gamma}^{-1}G^{\top})d_{y}=GK_{\gamma}^{-1}(\bar{r}_{1}+\gamma G^{\top}\bar{r}_{2})-\bar{r}_{2}\;.$ (12)
  3. 3.
    Solve the system $K_{\gamma}d_{x}=\bar{r}_{1}+\gamma G^{\top}\bar{r}_{2}-G^{\top}d_{y}$ to recover the primal descent direction.
The method employs a sparse Cholesky factorization along with the conjugate gradient (Cg) algorithm hestenes-stiefel-1952 . The sparse Cholesky factorization is advantageous as it is stable without numerical pivoting, making the algorithm tractable on a GPU. Each Cg iteration requires applying sparse triangular solves with the factors of $K_{\gamma}$. For this reason, HyKKT is efficient only if the Cg solver converges in few iterations. Fortunately, the eigenvalues of the Schur complement $S_{\gamma}:=GK_{\gamma}^{-1}G^{\top}$ all converge to $\frac{1}{\gamma}$ as we increase the regularization parameter $\gamma$ (regev2023hykkt, , Theorem 4), implying that $\lim_{\gamma\to\infty}\kappa_{2}(S_{\gamma})=1$. Since the convergence of the Cg method depends on the number of distinct eigenvalues of $S_{\gamma}$, the larger the $\gamma$, the faster the convergence of the Cg algorithm in (12). Although we observe similar performance, these methods ensure a monotonic decrease in the residual norm of (12) at each iteration.
Remark 2
The parameter $\gamma$ should be set above a given threshold $\underline{\gamma}$ for $K_{\gamma}$ to ensure positive definiteness of the system in (12). However, from the numerical stand-point, it is desired to choose moderate values of $\gamma$, since large $\gamma$ increases the ill-conditioning in $K_{\gamma}$, hence deteriorating the quality of the solution. In golub2003solving , the authors recommend choosing $\gamma=\|K\|/\|G\|^{2}$, which offers a good empirical trade-off. In (gould2010spectral, , Lemma 2.10), a tighter estimate is provided for the threshold $\underline{\gamma}$ if $G$ has full row-rank and $Z^{\top}KZ$ is positive definite:
$\underline{\gamma}=\dfrac{1}{\sigma_{\min}^{2}}\left(\dfrac{\|K\|^{2}}{\lambda_{\min}^{Z}}-\lambda_{\min}^{Y}\right)\;,$ (13)
with $\sigma_{\min}$ being the smallest nonzero singular value of the Jacobian $G$, $\lambda_{\min}^{Z}>0$ the smallest eigenvalue of $Z^{\top}KZ$, and $\lambda_{\min}^{Y}$ the smallest eigenvalue of $Y^{\top}KY$. From (13), we observe that $\underline{\gamma}$ is inversely proportional to $\lambda_{\min}^{Z}$ and $\sigma_{\min}$, indicating that it can reach a large magnitude if $G$ has nearly aligned rows or if $Z^{\top}KZ$ is close to being indefinite. In other words, HyKKT may have difficulties in solving problems with nearly dependent equality constraints or when the reduced Hessian is close to being indefinite.
3.2 Equality Relaxation Strategy: LiftedKKT
For a small relaxation parameter $\tau>0$ (chosen based on the numerical tolerance of the optimization solver $\varepsilon_{tol}$), the equality relaxation strategy shin2023accelerating approximates the equalities with lifted inequalities:
$\min_{x\in\mathbb{R}^{n}}\;f(x)\quad\text{subject to}\quad-\tau\leq g(x)\leq\tau\;,~h(x)\leq 0\;.$ (14)
The problem (14) has only inequality constraints. As a result, the condensed KKT system ($K_{1}$) reduces to
$K_{\tau}\,d_{x}=-r_{1}-H_{\tau}^{\top}(D_{H}r_{4}-Cr_{2})\;,$ (15)
with $H_{\tau}=\big{(}G^{\top}~H^{\top}\big{)}^{\top}$ and $K_{\tau}:=W+\delta_{w}I+H_{\tau}^{\top}D_{H}H_{\tau}$ (recall $\delta_{c}=0$ is assumed per Remark 1) . Using the relation (9), the matrix $K_{\tau}$ is guaranteed to be positive definite if the primal regularization parameter $\delta_{w}$ is adequately large. As such, the parameter $\delta_{w}$ is chosen dynamically using the inertia information of the system in ($K_{1}$). Therefore, $K_{\tau}$ can be factorized with a Cholesky decomposition, satisfying the key requirement of stable pivoting for implementation on the GPU. The relaxation introduces an error in the final solution. Fortunately, the error is of the same order as the solver tolerance, thus it does not significantly deteriorate the solution quality for small $\varepsilon_{tol}$.
While this method can be implemented with minor modifications in the optimization solver, the presence of tight inequalities in (14) causes severe ill-conditioning throughout the IPM iterations. Therefore, using an accurate iterative refinement algorithm is necessary to achieve reliable convergence behavior.
3.3 Discussion
We have introduced two algorithms to solve KKT systems on the GPU. Unlike classical implementations, the two methods do not require computing a sparse $\mathrm{LBL^{T}}$ factorization of the KKT system and instead utilize alternate reformulations based on the condensed KKT system ($K_{1}$). Both strategies rely on a Cholesky factorization: HyKKT factors a positive definite matrix $K_{\gamma}$ obtained using an Augmented Lagrangian strategy, whereas LiftedKKT factors a positive definite matrix $K_{\tau}$ after employing an equality relaxation strategy. We will demonstrate in the next section that the ill-conditioned matrices $K_{\gamma}$ and $K_{\tau}$ possess a specific structure that limits the loss of accuracy in IPM.
4 Conditioning of the condensed KKT system
The condensed matrix $K$ appearing in ($K_{1}$) is known to be increasingly ill-conditioned as the primal-dual iterates approach to a local solution with active inequalities. This behavior is amplified for the matrices $K_{\gamma}$ and $K_{\tau}$, as HyKKT and LiftedKKT require the use of respectively large $\gamma$ and small $\tau$. In this section, we analyze the numerical error associated with the solution of the condensed KKT system and discuss how the structured ill-conditioning makes the adverse effect of ill-conditioning relatively benign.
We first discuss the perturbation bound for a generic linear system $Mx=b$. The relative error after perturbing the right-hand side by $\Delta b$ is bounded by:
$\|\Delta x\|\leq\|M^{-1}\|\|\Delta b\|\;,\quad\frac{\|\Delta x\|}{\|x\|}\leq\kappa_{2}(M)\,\frac{\|\Delta b\|}{\|b\|}\;.$ (16a)
When the matrix is perturbed by $\Delta M$, the perturbed solution $\widehat{x}$ satisfies $\Delta x=\widehat{x}-x=-(M+\Delta M)^{-1}\Delta M\widehat{x}$. If $\kappa_{2}(M)\approx\kappa_{2}(M+\Delta M)$, we have $M\Delta x\approx-\Delta Mx$ (neglecting second-order terms), giving the bounds
$\|\Delta x\|\lessapprox\|M^{-1}\|\|\Delta M\|\|x\|\;,\quad\frac{\|\Delta x\|}{\|x\|}\lessapprox\kappa_{2}(M)\frac{\|\Delta M\|}{\|M\|}\;.$ (16b)
The relative errors are bounded above by a term depending on the conditioning $\kappa_{2}(M)$. Hence, it is legitimate to investigate the impact of the ill-conditioning when solving the condensed system ($K_{1}$) with LiftedKKT or with HyKKT. We will see that we can tighten the bounds in (16) by exploiting the structured ill-conditioning of the condensed matrix $K$. We base our analysis on wright1998ill , where the author has put a particular emphasis on the condensed KKT system ($K_{1}$) without equality constraints. We generalize their results to the matrix $K_{\gamma}$, which incorporates both equality and inequality constraints. The results extend directly to $K_{\tau}$ (by letting the number of equalities to be zero).
4.1 Centrality conditions
We start the discussion by recalling important results about the iterates of the interior-point algorithm. For $p:=(x,s,y,z)$, we denote by $(p,v)$ the current primal-dual iterate, and $(p^{\star},v^{\star})$ a solution of the KKT conditions (4). We denote $\mathcal{B}=\mathcal{B}(x^{\star})$ the active-set at the optimal solution $x^{\star}$, and $\mathcal{N}=\mathcal{N}(x^{\star})$ the inactive set. In this section, we are interested in the local convergence behavior of the primal-dual iterate, and we suppose $(p,v)$ is close enough to the solution $(p^{\star},v^{\star})$.
Assumption 4.1
Let $(p^{\star},v^{\star})$ be a primal-dual solution satisfying the KKT conditions (4). Let the following hold:
  • Continuity: The Hessian $\nabla^{2}_{xx}L(\cdot)$ is Lipschitz continuous near $p^{\star}$;
  • Linear Independence Constraint Qualification (LICQ): the active Jacobian $A(x^{\star})$ is full row-rank;
  • Strict Complementarity (SCS): for every $i\in\mathcal{B}(x^{\star})$, $z_{i}^{\star}>0$.
  • Second-order sufficiency (SOSC): for every $h\in\operatorname{Ker}\!\big{(}A(x^{\star})\big{)}$, $h^{\top}\nabla_{xx}^{2}L(p^{\star})h>0$.
We denote $\delta(p,v)=\|(p,v)-(p^{\star},v^{\star})\|$ the Euclidean distance to the primal-dual stationary point $(p^{\star},v^{\star})$. From (wright2001effects, , Theorem 2.2), if Assumption 4.1 holds at $p^{\star}$ and $v>0$,
[…]
For feasible iterate $(s,v)>0$, we define the duality measure $\Xi(s,v)$ as the mapping
$\Xi(s,v)=s^{\top}v/m_{i}\;,$ (18)
where $m_{i}$ is the number of inequality constraints. The duality measure encodes the current satisfaction of the complementarity constraints. For a solution $(p^{\star},v^{\star})$, we have $\Xi(s^{\star},v^{\star})=0$. The duality measure can be used to define the barrier parameter in IPM.
We suppose the iterates $(p,v)$ satisfy the centrality conditions
$\|\nabla_{p}\mathcal{L}(p,v)\|\leq C\;\Xi(s,v)\;,$ (19a)
$(s,v)>0\;,\quad s_{i}v_{i}\geq\alpha\,\Xi(s,v)\quad\forall i=1,\cdots,m_{i}\;,$ (19b)
for some constants $C>0$ and $\alpha\in(0,1)$. Conditions (19b) ensure that the products $s_{i}v_{i}$ are not too disparate in the diagonal term $D_{s}$. This condition is satisfied (even if rather loosely) in the solver Ipopt (see (wachter2006implementation, , Equation (16))).
Proposition 1 (wright2001effects , Lemma 3.2 and Theorem 3.3)
Suppose $p^{\star}$ satisfies Assumption 4.1. If the current primal-dual iterate $(p,v)$ satisfies the centrality conditions (19), then
$i\in\mathcal{B}\implies s_{i}=\Theta(\Xi)\,,\quad v_{i}=\Theta(1)\;,$ (20a)
$i\in\mathcal{N}\implies s_{i}=\Theta(1)\,,\quad v_{i}=\Theta(\Xi)\;.$ (20b)
and the distance to the solution $\delta(p,v)$ is bounded by the duality measure $\Xi$:
$\delta(p,v)=O(\Xi)\;.$ (20c)
4.2 Structured ill-conditioning
The following subsection looks at the structure of the condensed matrix $K_{\gamma}$ in HyKKT. All the results apply directly to the matrix $K_{\tau}$ in LiftedKKT, by setting the number of equality constraints to $m_{e}=0$. First, we show that if the iterates $(p,v)$ satisfy the centrality conditions (19), then the condensed matrix $K_{\gamma}$ exhibits a structured ill-conditioning.
4.2.1 Invariant subspaces in $K_{\gamma}$
Without regularization we have that $K_{\gamma}=W+H^{\top}D_{s}H+\gamma G^{\top}G$, with the diagonal $D_{s}=S^{-1}V$. We note by $m_{a}$ the cardinality of the active set $\mathcal{B}$, $H_{\mathcal{B}}$ the Jacobian of active inequality constraints, $H_{\mathcal{N}}$ the Jacobian of inactive inequality constraints and by $A:=\begin{bmatrix}H_{\mathcal{B}}^{\top}&G^{\top}\end{bmatrix}^{\top}$ the active Jacobian. We define the minimum and maximum active slack values as
$s_{min}=\min_{i\in\mathcal{B}}s_{i}\;,\quad s_{max}=\max_{i\in\mathcal{B}}s_{i}\;.$ (21)
We recall that $m_{e}$ is the number of equality constraints, and define $\ell:=m_{e}+m_{a}$.
We express the structured ill-conditioning of $K_{\gamma}$ by modifying the approach outlined in (wright1998ill, , Theorem 3.2) to account for the additional term $\gamma G^{\top}G$ arising from the equality constraints. We show that the matrix $K_{\gamma}$ has two invariant subspaces (in the sense defined in (stewart1990matrix, , Chapter 5)), associated respectively to the range of the transposed active Jacobian (large space) and to the null space of the active Jacobian (small space).
Theorem 4.2 (Properties of $K_{\gamma}$)
Suppose the condensed matrix is evaluated at a primal-dual point $(p,\nu)$ satisfying (19), for sufficiently small $\Xi$. Let $\lambda_{1},\cdots,\lambda_{n}$ be the $n$ eigenvalues of $K_{\gamma}$, ordered as $|\lambda_{1}|\geq\cdots\geq|\lambda_{n}|$. Let $\begin{bmatrix}Y&Z\end{bmatrix}$ be an orthogonal matrix, where $Z$ encodes the basis of the null-space of $A$. Let $\underline{\sigma}:=\min\left(\frac{1}{\Xi},\gamma\right)$ and $\overline{\sigma}:=\max\left(\frac{1}{s_{min}},\gamma\right)$. Then,
  1. (i)
    The $\ell$ largest-magnitude eigenvalues of $K_{\gamma}$ are positive, with $\lambda_{1}=\Theta(\overline{\sigma})$ and $\lambda_{\ell}=\Omega(\underline{\sigma})$.
  2. (ii)
    The $n$-$\ell$ smallest-magnitude eigenvalues of $K_{\gamma}$ are $\Theta(1)$.
  3. (iii)
    If $0<\ell<n$, then $\kappa_{2}(K_{\gamma})=\Theta(\overline{\sigma})$.
  4. (iv)
    There are orthonormal matrices $\widetilde{Y}$ and $\widetilde{Z}$ for simple invariant subspaces of $K_{\gamma}$ such that $Y-\widetilde{Y}=O(\underline{\sigma}^{-1})$ and $Z-\widetilde{Z}=O(\underline{\sigma}^{-1})$.
Proof
We start the proof by setting apart the inactive constraints from the active constraints in $K_{\gamma}$:
$K_{\gamma}=W+H_{\mathcal{N}}^{\top}S_{\mathcal{N}}^{-1}V_{\mathcal{N}}H_{\mathcal{N}}+A^{\top}D_{\gamma}A\,,\quad\text{with}\quad D_{\gamma}=\begin{bmatrix}S_{\mathcal{B}}^{-1}V_{\mathcal{B}}&0\\ 0&\gamma I\end{bmatrix}\;$. (22)
Using Assumption 4.1, Lipschitz continuity implies that the Hessian and the inactive Jacobian are bounded: $W=O(1)$, $H_{\mathcal{N}}=O(1)$. Proposition 1 implies that $s_{\mathcal{N}}=\Theta(1)$ and $v_{\mathcal{N}}=\Theta(\Xi)$. We deduce:
$H_{\mathcal{N}}^{\top}S_{\mathcal{N}}^{-1}V_{\mathcal{N}}H_{\mathcal{N}}=O(\Xi)\;.$ (23)
Hence, for small enough $\Xi$, the condensed matrix $K_{\gamma}$ is dominated by the block of active constraints:
$K_{\gamma}=A^{\top}D_{\gamma}A+O(1)\;.$ (24)
Sufficiently close to the optimum $p^{\star}$, the constraints qualification in Assumption 4.1 implies that $A=\Theta(1)$ and has rank $\ell$. The eigenvalues $\{\eta_{i}\}_{i=1,\cdots,n}$ of $A^{\top}D_{\gamma}A$ satisfy $\eta_{i}>0$ for $i=1,\cdots,\ell$ and $\eta_{i}=0$ for $i=\ell+1,\cdots,n$. As $s_{\mathcal{B}}=\Theta(\Xi)$ and $v_{\mathcal{B}}=\Theta(1)$ (Proposition 1), the smallest diagonal element in $D_{\gamma}$ is $\Omega(\min\{\frac{1}{\Xi},\gamma\})$ and the largest diagonal element is $\Theta(\max\{\frac{1}{s_{min}},\gamma\})$. Hence,
$\eta_{1}=\Theta(\overline{\sigma})\;,\quad\eta_{\ell}=\Omega(\underline{\sigma})\;.$ (25)
Using (wright1998ill, , Lemma 3.1), we deduce $\lambda_{1}=\Theta(\overline{\sigma})$ and $\lambda_{\ell}=\Omega(\underline{\sigma})$, proving the first result (i).
Let $L_{\gamma}:=A^{\top}D_{\gamma}A$. We have that
$\begin{bmatrix}Z^{\top}\\ Y^{\top}\end{bmatrix}L_{\gamma}\begin{bmatrix}Z&Y\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ 0&L_{2}\end{bmatrix}\;$, (26)
with $L_{1}=0$ and $L_{2}=Y^{\top}L_{\gamma}Y$. The smallest eigenvalue of $L_{2}$ is $\Omega(\underline{\sigma})$ and the matrix $E:=K_{\gamma}-L_{\gamma}$ is $O(1)$. By applying (wright1998ill, , Theorem 3.1, (ii)), the $n$-$\ell$ smallest eigenvalues in $K_{\gamma}$ differ by $\Omega(\underline{\sigma}^{-1})$ from those of the reduced Hessian $Z^{\top}K_{\gamma}Z$. In addition, (23) implies that $Z^{\top}K_{\gamma}Z-Z^{\top}WZ=O(\Xi)$. Using SOSC, $Z^{\top}WZ$ is positive definite for small enough $\Xi$, implying all its eigenvalues are $\Theta(1)$. Using again (wright1998ill, , Lemma 3.1), we get that the $n$-$\ell$ smallest eigenvalues in $K_{\gamma}$ are $\Theta(1)$, proving (ii). The results in (iii) can be obtained by combining (i) and (ii) (provided $0<\ell<n$). Finally, point (iv) directly follows from (wright1998ill, , Theorem 3.1 (i)).
Corollary 1
The condensed matrix $K_{\gamma}$ can be decomposed as
$K_{\gamma}=U\Sigma U^{\top}=\begin{bmatrix}U_{L}&U_{S}\end{bmatrix}\begin{bmatrix}\Sigma_{L}&0\\ 0&\Sigma_{S}\end{bmatrix}\begin{bmatrix}U_{L}^{\top}\\ U_{S}^{\top}\end{bmatrix}\;$, (27)
with $\Sigma_{L}=\operatorname{diag}(\sigma_{1},\cdots,\sigma_{\ell})\in\mathbb{R}^{\ell\times\ell}$ and $\Sigma_{S}=\operatorname{diag}(\sigma_{\ell+1},\cdots,\sigma_{n})\in\mathbb{R}^{(n-\ell)\times(n-\ell)}$ two diagonal matrices, and $U_{L}\in\mathbb{R}^{n\times\ell}$, $U_{S}\in\mathbb{R}^{n\times(n-\ell)}$ two orthogonal matrices such that $U_{L}^{\top}U_{S}=0$. The diagonal elements in $\Sigma_{S}$ and $\Sigma_{L}$ satisfy
$\frac{\sigma_{1}}{\sigma_{\ell}}\ll\frac{\sigma_{1}}{\sigma_{n}}\;,\quad\frac{\sigma_{\ell+1}}{\sigma_{n}}\ll\frac{\sigma_{1}}{\sigma_{n}}\;.$ (28)
For suitably chosen basis $Y$ and $Z$, spanning respectively the row space and the null space of the active Jacobian $A$, we get
$U_{L}-Y=O(\underline{\sigma}^{-1})\;,\quad U_{S}-Z=O(\underline{\sigma}^{-1})\;.$ (29)
Proof
Using the spectral theorem, we obtain the decomposition as (27). According to Theorem 4.2, the $\ell$ largest eigenvalues of $K_{\gamma}$ are large and well separated from the $n$-$\ell$ smallest eigenvalues, establishing (28). Using Theorem 4.2, part (iv), we obtain the result in (29).
Corollary 1 gives us a deeper insight into the structure of the condensed matrix $K_{\gamma}$. Using equation (29), we observe we can assimilate the large space of $K_{\gamma}$ with $\operatorname{range}(A^{\top})$ and the small space with $\operatorname{Ker}(A)$. The decomposition (27) leads to the following relations
$\|K_{\gamma}\|=\|\Sigma_{L}\|=\Theta(\overline{\sigma})\;,$ $\Sigma_{L}^{-1}=O(\underline{\sigma}^{-1})\;,$ (30)
$\|K_{\gamma}^{-1}\|=\|\Sigma_{S}^{-1}\|=\Theta(1)\,$, $\Sigma_{S}=\Theta(1)\,.$
The condition of $\Sigma_{L}$ depends on $\kappa_{2}(A)$ and the ratio $\frac{s_{max}}{s_{min}}=O(\Xi\overline{\sigma})$. The condition of $\Sigma_{S}$ reflects the condition of the reduced Hessian $Z^{\top}WZ$.
Three observations are due:
  1. 1.
    Theorem 4.2 (iii) tells us that $\kappa_{2}(K_{\gamma})=\Theta(\overline{\sigma})$, meaning that if $\gamma\geq\frac{1}{s_{min}}$, then the conditioning $\kappa_{2}(K_{\gamma})$ increases linearly with $\gamma$, hence recovering a known result regev2023hykkt .
  2. 2.
    In early IPM iterations, the slacks are pushed away from the boundary and the number of active inequality constraints is $m_{a}=0$. The ill-conditioning in $K_{\gamma}$ is caused only by $\gamma G^{\top}G$ and $\underline{\sigma}=\overline{\sigma}=\gamma$.
  3. 3.
    In late IPM iterations, the active slacks are converging to 0. We observe that if $\frac{1}{\Xi}\leq\gamma\leq\frac{1}{s_{min}}$ the parameter $\gamma$ does not increase the ill-conditioning of the condensed matrix $K_{\gamma}$.
4.2.2 Numerical accuracy of the condensed matrix $K_{\gamma}$
In floating-point arithmetic, the condensed matrix $K_{\gamma}$ is evaluated as
$\widehat{K}_{\gamma}=W+\Delta W+(A+\Delta A)^{\top}(D_{\gamma}+\Delta D_{\gamma})(A+\Delta A)\\ +(H_{\mathcal{N}}+\Delta H_{\mathcal{N}})^{\top}S_{\mathcal{N}}^{-1}V_{\mathcal{N}}(H_{\mathcal{N}}+\Delta H_{\mathcal{N}})\;,$
with $\Delta W=O(\mathbf{u})$, $\Delta H_{\mathcal{N}}=O(\mathbf{u})$, $\Delta A=\Theta(\mathbf{u})$, $\Delta D_{\gamma}=O(\mathbf{u}\overline{\sigma})$: most of the errors arise because of the ill-conditioned diagonal terms in $D_{\gamma}$.
Proposition 2
In floating-point arithmetic, the perturbation of the condensed matrix $K_{\gamma}$ satisfies $\Delta K_{\gamma}:=\widehat{K_{\gamma}}-K_{\gamma}=O(\mathbf{u}\overline{\sigma})$.
Proof
As $A=\Theta(1)$, we have: $A^{\top}D_{\gamma}A=\Theta(\overline{\sigma})$ and $A^{\top}\Delta D_{\gamma}A=O(\mathbf{u}\overline{\sigma})$. Neglecting second-order terms, we get
$\Delta K_{\gamma}=\overbrace{\Delta W}^{O(\mathbf{u})}+\overbrace{\Delta A^{\top}D_{\gamma}A}^{O(\overline{\sigma}\mathbf{u})}+\overbrace{A^{\top}D_{\gamma}\Delta A}^{O(\overline{\sigma}\mathbf{u})}+\overbrace{A^{\top}\Delta D_{\gamma}A}^{O(\overline{\sigma}\mathbf{u})}\\ +\underbrace{\Delta H_{\mathcal{N}}S_{\mathcal{N}}^{-1}V_{\mathcal{N}}H_{\mathcal{N}}}_{O(\mathbf{u})}+\underbrace{H_{\mathcal{N}}S_{\mathcal{N}}^{-1}V_{\mathcal{N}}\Delta H_{\mathcal{N}}}_{O(\mathbf{u})}\;,$
where the terms in braces show the respective bounds on the errors. We deduce the error is dominated by the terms arising from the active Jacobian, all bounded by $O(\overline{\sigma}\mathbf{u})$, hence concluding the proof.
If it is large enough, the unstructured perturbation $\Delta K_{\gamma}$ can impact the structured ill-conditioning in the perturbed matrix $\widehat{K}_{\gamma}$. We know that the smallest eigenvalue $\eta_{\ell}$ of $A^{\top}D_{\gamma}A$ is $\Omega(\underline{\sigma})$. As mentioned in (wright1998ill, , Section 3.4.2), the perturbed matrix $\widehat{K}_{\gamma}$ keeps the $p$ large eigenvalues bounded below by $\underline{\sigma}$ if the perturbation is itself much smaller than the eigenvalue $\eta_{\ell}$:
$\|\Delta K_{\gamma}\|\ll\eta_{\ell}=\Omega(\underline{\sigma})\;.$ (31)
However, the bound given in Proposition 2 is too loose for (31) to hold without any further assumption (we have only $\underline{\sigma}\leq\overline{\sigma}$). We note that for some constant $C>0$, $\Delta K_{\gamma}\leq C\mathbf{u}\overline{\sigma}$, implying $\Delta K_{\gamma}/\underline{\sigma}\leq C\mathbf{u}\overline{\sigma}/\underline{\sigma}$. Hence, if we suppose in addition the ratio $\overline{\sigma}/\underline{\sigma}$ is close to $1$, then $\|\Delta K_{\gamma}\|=O(\mathbf{u}\overline{\sigma})$ can instead be replaced by $\|\Delta K_{\gamma}\|=O(\mathbf{u}\underline{\sigma})$, ensuring (31) holds.
4.2.3 Numerical solution of the condensed system
We are interested in estimating the relative error made when solving the system $K_{\gamma}x=b$ in floating point arithmetic. We suppose $K_{\gamma}$ is factorized using a backward-stable Cholesky decomposition. The computed solution $\widehat{x}$ is solution of a perturbed system $\widetilde{K}_{\gamma}\widehat{x}=b$, with $\widetilde{K}_{\gamma}=K_{\gamma}+\Delta_{s}K_{\gamma}$ and $\Delta_{s}K_{\gamma}$ a symmetric matrix satisfying
$\|\Delta_{s}K_{\gamma}\|\leq\mathbf{u}\varepsilon_{n}\|K_{\gamma}\|\;,$ (32)
for $\varepsilon_{n}$ a small constant depending on the dimension $n$. We need the following additional assumptions to ensure (a) the Cholesky factorization runs to completion and (b) we can incorporate the backward-stable perturbation $\Delta_{s}K_{\gamma}$ in the generic perturbation $\Delta K_{\gamma}$ introduced in Proposition 2.
Assumption 4.3
Let $(p,v)$ be the current primal-dual iterate. We assume:
  • (a)
    $(p,v)$ satisfies the centrality conditions (19).
  • (b)
    The parameter $\gamma$ satisfies $\gamma=\Theta(\Xi^{-1})$.
  • (c)
    The duality measure is large enough relative to the precision $\mathbf{u}$: $\mathbf{u}\ll\Xi$.
  • (d)
    The primal step $\widehat{x}$ is computed using a backward stable method satisfying (32) for a small constant $\varepsilon_{n}$.
Condition (a) implies both $s_{min}=\Theta(\Xi)$ and $s_{max}=\Theta(\Xi)$ (Proposition 1). Condition (b) supposes in addition $\gamma=\Theta(\Xi^{-1})$, making the matrix $\Sigma_{L}$ well-conditioned with $\underline{\sigma}=\Theta(\Xi^{-1})$, $\overline{\sigma}=\Theta(\Xi^{-1})$ and $\overline{\sigma}/\underline{\sigma}=\Theta(1)$. This is in line with the recommendation in golub2003solving , where the authors recommend chosing $\gamma=\frac{\|K\|}{\|G\|^{2}}$ (as here the condition (a) implies that $\|K\|=\Theta(\Xi)$). Condition (c) ensures that $\kappa_{2}(K_{\gamma})=\Theta(\overline{\sigma})$ satisfies $\kappa_{2}(K_{\gamma})\mathbf{u}\ll 1$ (implying the Cholesky factorization runs to completion). Condition (d) tells us that the perturbation caused by the Cholesky factorization is $\Delta_{s}K_{\gamma}=O(\mathbf{u}\|K_{\gamma}\|)$. As (30) implies $\|K_{\gamma}\|=\Theta(\Xi^{-1})$, we can incorporate $\Delta_{s}K_{\gamma}$ in the perturbation $\Delta K_{\gamma}$ given in Proposition 2.
We are now ready to analyze the perturbation bound for the condensed system. We denote $x$ the solution of the linear system $K_{\gamma}x=b$ in exact arithmetic, and $\widehat{x}$ the solution of the perturbed system $\widehat{K}_{\gamma}\widehat{x}=\widehat{b}$ in floating-point arithmetic. We are interested in bounding the error $\Delta x=\widehat{x}-x$. We recall that every vector $x\in\mathbb{R}^{n}$ decomposes as
$x=U_{L}x_{L}+U_{S}x_{S}=Yx_{Y}+Zx_{Z}\;.$ (33)
Impact of right-hand-side perturbation.
Using (27), the inverse of $K_{\gamma}$ satisfies
$K_{\gamma}^{-1}=\begin{bmatrix}U_{L}&U_{S}\end{bmatrix}\begin{bmatrix}\Sigma_{L}^{-1}&0\\ 0&\Sigma_{S}^{-1}\end{bmatrix}\begin{bmatrix}U_{L}^{\top}\\ U_{S}^{\top}\end{bmatrix}\;$. (34)
Hence, if we solve the system for $\widehat{b}:=b+\Delta b$, $\Delta x=K_{\gamma}^{-1}\Delta b$ decomposes as
$\begin{bmatrix}\Delta x_{L}\\ \Delta x_{S}\end{bmatrix}=\begin{bmatrix}\Sigma_{L}^{-1}&0\\ 0&\Sigma_{S}^{-1}\end{bmatrix}\begin{bmatrix}\Delta b_{L}\\ \Delta b_{S}\end{bmatrix}\;$, (35)
which in turn implies the following bounds:
$\|\Delta x_{L}\|\leq\|\Sigma_{L}^{-1}\|\|\Delta b_{L}\|\;,\quad\|\Delta x_{S}\|\leq\|\Sigma_{S}^{-1}\|\|\Delta b_{S}\|\;.$ (36)
As $\Sigma_{L}^{-1}=O(\Xi)$ and $\Sigma_{S}^{-1}=\Theta(1)$, we deduce that the error $\Delta x_{L}$ is smaller by a factor of $\Xi$ than the error $\Delta x_{S}$. The total error $\Delta x=U_{L}\Delta x_{L}+U_{S}\Delta x_{S}$ is bounded by
$\|\Delta x\|\leq\|\Sigma_{L}^{-1}\|\|\Delta b_{L}\|+\|\Sigma_{S}^{-1}\|\|\Delta b_{S}\|=O(\|\Delta b\|)\;.$ (37)
Impact of matrix perturbation.
As $\|\Delta K_{\gamma}\|\ll\|K_{\gamma}\|$, we have that
$(K_{\gamma}+\Delta K_{\gamma})^{-1}$ $=(I+K_{\gamma}^{-1}\Delta K_{\gamma})^{-1}K_{\gamma}^{-1}\;,$ (38)
$=K_{\gamma}^{-1}-K_{\gamma}^{-1}\Delta K_{\gamma}K_{\gamma}^{-1}+O(\|\Delta K_{\gamma}\|^{2})\;.$
We decompose $\Delta K_{\gamma}$ in two matrices $\Gamma_{L}\in\mathbb{R}^{\ell\times n}$ and $\Gamma_{S}\in\mathbb{R}^{(n-\ell)\times n}$ such that $\Delta K_{\gamma}=\begin{bmatrix}\Gamma_{L}\\ \Gamma_{S}\end{bmatrix}$. Using (34) the first-order error is given by
$K_{\gamma}^{-1}\Delta K_{\gamma}K_{\gamma}^{-1}=U_{L}\Sigma_{L}^{-1}\Gamma_{L}\Sigma_{L}^{-1}U_{L}^{\top}+U_{S}\Sigma_{S}^{-1}\Gamma_{S}\Sigma_{S}^{-1}U_{S}^{\top}\;.$ (39)
Using (30) and $(\Gamma_{L},\Gamma_{S})=O(\Xi^{-1}\mathbf{u})$, we obtain $\Sigma_{L}^{-1}\Gamma_{L}\Sigma_{L}^{-1}=O(\Xi\mathbf{u})$ and $\Sigma_{S}^{-1}\Gamma_{S}\Sigma_{S}^{-1}=O(\Xi^{-1}\mathbf{u})$. We deduce that the error made in the large space is $O(\Xi\mathbf{u})$ whereas the error in the small space is $O(\Xi^{-1}\mathbf{u})$.
4.3 Solution of the condensed KKT system
We use the relations (36) and (39) to bound the error made when solving the condensed KKT system ($K_{1}$) in floating-point arithmetic. In all this section, we assume that the primal-dual iterate $(p,v)$ satisfies Assumption 4.3. Using (wright2001effects, , Corollary 3.3), the solution $(d_{x},d_{y})$ of the condensed KKT system ($K_{1}$) in exact arithmetic satisfies $(d_{x},d_{y})=O(\Xi)$. In ($K_{1}$), the RHS $\bar{r}_{1}$ and $\bar{r}_{2}$ evaluate in floating-point arithmetic as
$\left\{\begin{aligned} \bar{r}_{1}&=-\widehat{r}_{1}+\widehat{H}^{\top}\big{(}\widehat{D}_{s}\widehat{r}_{4}-\widehat{r}_{2}\big{)}\;,\\ \bar{r}_{2}&=-\widehat{r}_{3}\;.\end{aligned}\right$. (40)
Using basic floating-point arithmetic, we get $\widehat{r}_{1}=r_{1}+O(\mathbf{u})$, $\widehat{r}_{3}=r_{3}+O(\mathbf{u})$, $\widehat{r}_{4}=r_{4}+O(\mathbf{u})$. The error in the right-hand-side $r_{2}$ is impacted by the term $\mu S^{-1}e$: under Assumption 4.3, it impacts differently the active and inactive components: $\widehat{r}_{2,\mathcal{B}}=r_{2,\mathcal{B}}+O(\mathbf{u})$ and $\widehat{r}_{2,\mathcal{N}}=r_{2,\mathcal{N}}+O(\Xi\mathbf{u})$. Similarly, the diagonal matrix $\widehat{D}_{s}$ retains full accuracy only w.r.t. the inactive components: $\widehat{D}_{s,\mathcal{B}}=D_{s,\mathcal{B}}+O(\Xi^{-1}\mathbf{u})$ and $\widehat{D}_{s,\mathcal{N}}=D_{s,\mathcal{N}}+O(\Xi\mathbf{u})$.
4.3.1 Solution with HyKKT
We analyze the accuracy achieved when we solve the condensed system ($K_{1}$) using HyKKT, and show that the error remains reasonable even for large values of the regularization parameter $\gamma$.
Initial right-hand-side.
Let $\widehat{s}_{\gamma}:=\bar{r}_{1}+\gamma\widehat{G}^{\top}\bar{r}_{2}$. The initial right-hand side in (12) is evaluated as $\widehat{r}_{\gamma}:=\widehat{G}\widehat{K}_{\gamma}^{-1}\widehat{s}_{\gamma}-\bar{r}_{2}$. The following proposition shows that despite an expression involving the inverse of the ill-conditioned condensed matrix $K_{\gamma}$, the error made in $r_{\gamma}$ is bounded only by the machine precision $\mathbf{u}$.
Proposition 3
In floating point arithmetic, the error in the right-hand-side $\Delta\widehat{r}_{\gamma}$ satisfies:
$\Delta\widehat{r}_{\gamma}=-\Delta\bar{r}_{2}+\widehat{G}\widehat{K}_{\gamma}^{-1}\Delta s_{\gamma}=O(\mathbf{u})\;.$ (41)
Proof
Using (40), we have
$\bar{r}_{1}+\gamma\widehat{G}^{\top}\bar{r}_{2}$ $=-\widehat{r}_{1}+\gamma\widehat{G}^{\top}\widehat{r}_{3}+\widehat{H}^{\top}\big{(}\widehat{D}_{s}\widehat{r}_{4}-\widehat{r}_{2}\big{)}$
$=-\underbrace{\widehat{r}_{1}}_{O(\mathbf{u})}+\underbrace{\widehat{H}_{\mathcal{N}}^{\top}\big{(}\widehat{D}_{s,\mathcal{N}}\widehat{r}_{4,\mathcal{N}}-\widehat{r}_{2,\mathcal{N}}\big{)}}_{O(\Xi\mathbf{u})}+\underbrace{\widehat{A}^{\top}\begin{bmatrix}\widehat{D}_{s,\mathcal{B}}\widehat{r}_{4,\mathcal{B}}-\widehat{r}_{2,\mathcal{B}}\\ \gamma\widehat{r}_{3}\end{bmatrix}}_{O(\Xi^{-1}\mathbf{u})}\;.$
The error decomposes as $\Delta s_{\gamma}=Y\Delta s_{Y}+Z\Delta s_{Z}=U_{L}\Delta s_{L}+U_{S}\Delta s_{S}$. We have $\Delta s_{Y}=O(\Xi^{-1}\mathbf{u})$ and $\Delta s_{Z}=O(\mathbf{u})$. Using (29), we deduce $\Delta s_{L}=U_{L}^{\top}\Delta s_{\gamma}=O(\Xi^{-1}\mathbf{u})$ and $\Delta s_{S}=U_{S}^{\top}\Delta s_{\gamma}=O(\mathbf{u})$. Using (30) and (34), the error in the large space $\Delta s_{L}$ annihilates in the backsolve:
$K_{\gamma}^{-1}\Delta s_{\gamma}=U_{L}\Sigma_{L}^{-1}\Delta s_{L}+U_{S}\Sigma_{S}^{-1}\Delta s_{S}=O(\mathbf{u})\;.$ (42)
Finally, using (38), we get
$\widehat{G}\widehat{K}_{\gamma}^{-1}\Delta s_{\gamma}\approx\widehat{G}(I-K_{\gamma}^{-1}\Delta K_{\gamma})K_{\gamma}^{-1}\Delta s_{\gamma}\;.$ (43)
Using (42), the first term is $\widehat{G}K_{\gamma}^{-1}\Delta s_{\gamma}=O(\mathbf{u})$. We have in addition
$GK_{\gamma}^{-1}\Delta K_{\gamma}(K_{\gamma}^{-1}\Delta s_{\gamma})=\big{[}GU_{L}\Sigma_{L}^{-1}\Gamma_{L}+GU_{S}\Sigma_{S}^{-1}\Gamma_{S}\big{]}(K_{\gamma}^{-1}\Delta s_{\gamma})\;.$ (44)
Using again (29): $GU_{L}=GY+O(\Xi)$ and $GU_{S}=O(\Xi)$. Hence $GU_{L}\Sigma_{L}^{-1}\Gamma_{L}=O(1)$ and $GU_{S}\Sigma_{S}^{-1}\Gamma_{S}=O(1)$. Using (37), we have $K_{\gamma}^{-1}\Delta G^{\top}=O(\mathbf{u})$, implying $\Delta GK_{\gamma}^{-1}\Delta K_{\gamma}(K_{\gamma}^{-1}\Delta s_{\gamma})=O(\Xi^{-1}\mathbf{u}^{2})$. Assumption 4.3 implies that $\Xi^{-1}\mathbf{u}^{2}\ll\mathbf{u}$, proving (41).
Schur-complement operator.
The solution of the system (12) involves the Schur complement $S_{\gamma}=GK_{\gamma}^{-1}G^{\top}$. We show that the Schur complement has a specific structure that limits the loss of accuracy in the conjugate gradient algorithm.
Proposition 4
Suppose the current primal-dual iterate $(p,v)$ satisfies Assumption 4.3. In exact arithmetic,
$S_{\gamma}=GY\,\Sigma_{L}^{-1}\,Y^{\top}G^{\top}+O(\Xi^{2})\;.$ (45)
Proof
Using (34), we have
$GK_{\gamma}^{-1}G^{\top}=GU_{L}\Sigma_{L}^{-1}U_{L}^{\top}G^{\top}+GU_{S}\Sigma_{S}^{-1}U_{S}^{\top}G^{\top}\;.$ (46)
Using (29), we have $GU_{L}=GY+O(\Xi)$, and $G=O(1)$, implying
$GU_{L}\Sigma_{L}^{-1}U_{L}^{\top}G^{\top}=GY\Sigma_{L}^{-1}Y^{\top}G^{\top}+O(\Xi^{2})\;.$ (47)
Using again (29), we have $GU_{S}=O(\Xi)$. Hence, $GU_{S}\Sigma_{S}^{-1}U_{S}^{\top}G^{\top}=O(\Xi^{2})$, concluding the proof.
We adapt the previous proposition to bound the error made when evaluating $\widehat{S}_{\gamma}$ in floating-point arithmetic.
Proposition 5
Suppose the current primal-dual iterate $(p,v)$ satisfies Assumption 4.3. In floating-point arithmetic,
$\widehat{S}_{\gamma}=S_{\gamma}+O(\mathbf{u})\;.$ (48)
Proof
We denote $\widehat{G}=G+\Delta G$ (with $\Delta G=O(\mathbf{u})$). Then
$\widehat{S}_{\gamma}$ $=\widehat{G}\widehat{K}_{\gamma}^{-1}\widehat{G}^{\top}\;,$ (49)
$\approx(G+\Delta G)\big{(}K_{\gamma}^{-1}-K_{\gamma}^{-1}\Delta K_{\gamma}K_{\gamma}^{-1}\big{)}(G+\Delta G)^{\top}\;,$
$\approx S_{\gamma}-G\big{(}K_{\gamma}^{-1}\Delta K_{\gamma}K_{\gamma}^{-1}\big{)}G^{\top}+K_{\gamma}^{-1}\Delta G^{\top}+\Delta GK_{\gamma}^{-1}\;.$
The second line is given by (38), the third by neglecting the second-order errors. Using (37), we get $K_{\gamma}^{-1}\Delta G^{\top}=O(\mathbf{u})$ and $\Delta GK_{\gamma}^{-1}=O(\mathbf{u})$. Using (39), we have
$G\big{(}K_{\gamma}^{-1}\Delta K_{\gamma}K_{\gamma}^{-1}\big{)}G^{\top}=GU_{L}\Sigma_{L}^{-1}\Gamma_{L}\Sigma_{L}^{-1}U_{L}^{\top}G^{\top}+GU_{S}\Sigma_{S}^{-1}\Gamma_{S}\Sigma_{S}^{-1}U_{S}^{\top}G^{\top}\;.$
Using (29), we have $GU_{S}=O(\Xi)$. As $\Sigma_{S}^{-1}=\Theta(1)$ and $\Gamma_{S}=O(\Xi^{-1}\mathbf{u})$, we get $GU_{S}\Sigma_{S}^{-1}\Gamma_{S}\Sigma_{S}^{-1}U_{S}^{\top}G^{\top}=O(\Xi\mathbf{u})$. Finally, as $\Sigma_{L}^{-1}=\Theta(\Xi)$ and $GU_{L}=GY+O(\Xi)$, we have
$GU_{L}\Sigma_{L}^{-1}\Gamma_{L}\Sigma_{L}^{-1}U_{L}^{\top}G^{\top}=GY\Sigma_{L}^{-1}\Gamma_{L}\Sigma_{L}^{-1}Y^{\top}G^{\top}+O(\Xi^{2}\mathbf{u})\;.$ (50)
We conclude the proof by using $GY\Sigma_{L}^{-1}\Gamma_{L}\Sigma_{L}^{-1}Y^{\top}G^{\top}=O(\Xi\mathbf{u})$.
The two error bounds (41) and (48) ensure that we can solve (12) using a conjugate gradient algorithm, as the errors remain limited in floating-point arithmetic.
4.3.2 Solution with Lifted KKT system
The equality relaxation strategy used in LiftedKKT removes the equality constraints from the optimization problems, simplifying the solution of the condensed KKT system to (15). The active Jacobian $A$ reduces to the active inequalities $A=H_{\mathcal{B}}$, and we recover the original setting presented in wright1998ill . Using the same arguments than in (4.3.1), the error in the right-hand-side is bounded by $O(\Xi^{-1}\mathbf{u})$ and is in the range space of the active Jacobian $A$. Using (34), we can show that the absolute error on $\widehat{d}_{x}$ is bounded by $O(\Xi\mathbf{u})$. That implies the descent direction $\widehat{d}_{x}$ retains full relative precision close to optimality. In other words, we can refine the solution returned by the Cholesky solver accurately using Richardson iterations.
4.3.3 Summary
Numerically, the primal-dual step $(\widehat{d}_{x},\widehat{d}_{y})$ is computed only with an (absolute) precision $\varepsilon_{K}$, greater than the machine precision $\mathbf{u}$ (for HyKKT, $\varepsilon_{K}$ is the absolute tolerance of the Cg algorithm, for LiftedKKT the absolute tolerance of the iterative refinement algorithm).
The errors $\widehat{d}_{x}-d_{x}=O(\varepsilon_{K})$ and $\widehat{d}_{y}-d_{y}=O(\varepsilon_{K})$ propagate further in $(\widehat{d}_{s},\widehat{d}_{z})$. According to ($K_{1}$), we have $\widehat{d}_{s}=-\widehat{r}_{4}-\widehat{H}\widehat{d}_{x}$. By continuity, $\widehat{H}=H+O(\mathbf{u})$ and $\widehat{r}_{4}=r_{4}+O(\mathbf{u})$, implying
$\widehat{d}_{s}=d_{s}+O(\varepsilon_{K})\;.$ (51)
Eventually, we obtain $\widehat{d}_{z}=-\widehat{r}_{2}-\widehat{D}_{s}\widehat{d}_{s}$, giving the following bounds for the errors in the inactive and active components:
$\widehat{d}_{z,\mathcal{B}}$ $=-\widehat{r}_{2,\mathcal{B}}-\widehat{D}_{s,\mathcal{B}}\widehat{d}_{s,\mathcal{B}}=d_{z,\mathcal{B}}+O(\varepsilon_{K}\Xi^{-1})\;,$ (52)
$\widehat{d}_{z,\mathcal{N}}$ $=-\widehat{r}_{2,\mathcal{N}}-\widehat{D}_{s,\mathcal{N}}\widehat{d}_{s,\mathcal{N}}=d_{z,\mathcal{N}}+O(\varepsilon_{K}\Xi)\;.$
Most of the error arises in the descent w.r.t. the active bound’s multipliers $\widehat{d}_{z,\mathcal{B}}$. The impact remains limited if we have only few active inequality constraints.
5 Implementation
All our implementations use the Julia language bezanson-edelman-karpinski-shah-2017 . We utilized our local workstation to generate the results on the CPU, which is equipped with an AMD EPYC 7443 processor (3.2GHz) with 24 processors. For the GPU results, we used an NVIDIA A100 GPU (with CUDA 12.3) on the Polaris testbed at Argonne National Laboratory  [^1] https://www.alcf.anl.gov/polaris.
IPM solver.
We have implemented the two condensed-space methods in our nonlinear IPM solver MadNLP shin2021graph . This implementation utilizes the abstraction AbstractKKTSystem in MadNLP to represent the various formulations of the KKT linear systems. MadNLP can push most of the IPM algorithm to the GPU. In particular, any operation that involves the manipulation of array entries is performed by GPU kernels without transferring data to the host memory. We refer to shin2023accelerating for a detailed description of the GPU implementation in MadNLP.
Evaluation of the nonlinear models.
For fast evaluation of the derivatives on the GPU, we use the ExaModels.jl modeling tool shin2023accelerating to instantiate the evaluation of the nonlinear programs on the GPU. ExaModels.jl harnesses the sparsity structure and provides custom derivative kernels for repetitive algebraic subexpressions of the constraints and objective functions to compute first and second-order derivatives on the GPU in parallel bischof1991exploiting ; enzyme2021 . This approach caters to the SIMD architecture of the GPU by assigning each expression to multiple threads responsible for computing derivatives for different values. ExaModels.jl is efficient at evaluating the problem’s derivatives and offers a speedup often greater than 100 compared to standard tools like AMPL or JuMP (see Figure 1). As a result, the time spent in the evaluation of the model becomes negligible compared to the time spent in the linear solver. ExaModels.jl compiles the derivative evaluation kernels specifically for the given problem, whereas JuMP and AMPL depend on a precompiled interpreter to evaluate the model’s derivatives. Conversely, ExaModels.jl necessitates compilation time to create the derivative evaluation kernels, which can become substantial for problems involving a large number of heterogeneous constraints. Within this paper, we have always excluded the compilation time from the reported timings.
Figure 1: Evolution of the time spent evaluating the problem’s derivatives (gradient, Jacobian, Hessian) as we increase the problem’s dimension nh. In this benchmark, we compare four different instances from the COPS benchmark dolan2002benchmarking : rocket, robot, pinene, and dirichlet.
Linear solvers.
We solve the KKT systems assembled within MadNLP using various sparse linear solvers, chosen based on the KKT formulation $($$K_{1}$), ($K_{2}$), ($K_{3}$)) and the device (CPU, GPU) being utilized. We utilize the following solvers:
  • HSL MA27: Implements the multifrontal $\mathrm{LBL^{T}}$ factorization on the CPU duff1983multifrontal .
  • HSL MA57: Implements the multifrontal $\mathrm{LBL^{T}}$ factorization on the CPU duff1983multifrontal , using multithreaded BLAS kernels in the elimination algorithm.
  • HSL MA86: Implements a fine-grained multicore parallel supernodal $\mathrm{LBL^{T}}$ factorization on the CPU.
  • Panua Pardiso: Implements shared-memory parallel supernodal Cholesky, $\mathrm{LDL^{T}}$ and $\mathrm{LBL^{T}}$ factorizations on the CPU.
  • CHOLMOD: Implements the Cholesky and $\mathrm{LDL^{T}}$ factorizations on the CPU (using the AMD ordering amestoy-david-duff-2004 by default). It factorizes the condensed matrices $K_{\gamma}$ and $K_{\tau}$ appearing respectively in (11) and in (15).
  • cuDSS: Implements Cholesky, $\mathrm{LDL^{T}}$ and $\mathrm{LU}$ decompositions on an NVIDIA GPU. We use the $\mathrm{LDL^{T}}$ factorization to factorize the ill-conditioned condensed matrices $K_{\gamma}$ on the GPU, (here, we have observed that $\mathrm{LDL^{T}}$ is more robust than the Cholesky factorization).
  • Krylov.jl: Contains the Cg method used in the Golub & Greif strategy to solve (12) on both CPU and GPU architectures.
CHOLMOD chen-davis-hager-rajamanickam-2008 is shipped with Julia. For the HSL linear solvers, we utilize libHSL fowkes-lister-montoison-orban-2024 with the Julia interface HSL.jl montoison-orban-hsl-2021 . HSL MA57 and CHOLMOD are both compiled with OpenBLAS, a multithreaded version of BLAS and LAPACK. The Julia package Krylov.jl montoison2023krylov contains a collection of Krylov methods with a polymorphic implementation that can be used on both CPU and GPU architectures.
6 Numerical results
First, we assess in §6.1 the performance of the two hybrid solvers, LiftedKKT and HyKKT. The GPU implementation is compared to state-of-the-art solvers running on the CPU. Then, we present in §6.2 the results reported on the PGLIB OPF benchmark, complemented in §6.3 by the CUTEst benchmark.
6.1 Performance analysis of HyKKT and LiftedKKT on a large-scale instance
We evaluate the performance of each KKT solver on a large-scale OPF instance, taken from the PGLIB benchmark babaeinejadsarookolaee2019power : 78484epigrids. Our formulation with ExaModels has a total of 674,562 variables, 661,017 equality constraints, and 378,045 inequality constraints. Our previous work has pointed out that as soon as the OPF model is evaluated on the GPU using ExaModels, the time spent in AD becomes negligible, and the KKT solver becomes the bottleneck in the numerical implementation shin2023accelerating .
6.1.1 Performance comparison of linear solvers on CPUs and GPUs
We evaluate the individual performance of the cuDSS solver when factorizing the matrix $K_{\gamma}$ at the first IPM iteration (here with $\gamma=10^{7}$). We compare the times to perform the symbolic analysis, the factorization, and the triangular solves with those reported in CHOLMOD (running single-threaded), HSL MA86, and Panua Pardiso with sparse Cholesky. The CPU implementations achieved their highest efficiency when configured with 8 threads. Therefore, in the remainder of this paper, we will consistently present the performance of the linear solvers operating in parallel using 8 threads.
The results are displayed in Table 1. We benchmark in parallel the three decompositions implemented in cuDSS (Cholesky, $\mathrm{LDL^{T}}$, $\mathrm{LU}$). We observe that the analysis phase is four times slower for cuDSS compared to CHOLMOD. The ordering happening within the symbolic analysis is performed on CPU, so it cannot benefit from the GPU parallelism nvidiaNVIDIACuDSSPreview . Fortunately, the analysis can be performed only once, and the symbolic factorization can be reused if the optimization problem is re-solved with the same sparsity structure. The numerical factorization is about two times faster in cuDSS than in HSL MA86 , with a time almost independent of the factorization being used.
analysis (s) factorization (s) backsolve (s)
CHOLMOD $1.37\times 10^{0}$ $7.44\times 10^{-1}$ $5.20\times 10^{-2}$
PARDISO $4.52\times 10^{0}$ $2.94\times 10^{-1}$ $8.94\times 10^{-2}$
MA86 $3.88\times 10^{0}$ $1.38\times 10^{-1}$ $5.50\times 10^{-2}$
cuDSS-cholesky $3.63\times 10^{0}$ $4.05\times 10^{-2}$ $7.22\times 10^{-3}$
cuDSS-ldl $3.57\times 10^{0}$ $4.21\times 10^{-2}$ $6.35\times 10^{-3}$
cuDSS-lu $3.55\times 10^{0}$ $6.28\times 10^{-2}$ $6.42\times 10^{-3}$
Table 1: Comparing the performance of cuDSS with CHOLMOD , Pardiso, and HSL MA86 (both running in parallel using 8 threads) . The column for “analysis” displays the time spent in the symbolic analysis (computation of a reordering to minimize fill-in and construction of the elimination tree). The columns for “factorization” and “backsolves” display the time spent in the numerical factorization and in the ensuing triangular solve, respectively. The matrix $K_{\gamma}$ is symmetric positive definite, with a size $n=674,562$. The matrix is extremely sparse, with only $7,342,680$ non-zero entries (0.0020.002%).
6.1.2 Tuning the Golub & Greif strategy
We observe that the higher the regularization $\gamma$, the faster the Cg algorithm. However, we have to pay a price in terms of accuracy. In Figure 2, we depict the evolution of the number of Cg iterations and relative accuracy as we increase the parameter $\gamma$ from $10^{4}$ to $10^{8}$ in HyKKT. We observe that the number of CG iterations decreases by a factor of 10 as we increase $\gamma$, but the norm of the relative residual increases linearly as we increase $\gamma$. Thus, a trade-off exists between the time spent in Cg and the accuracy of the solution.
On the numerical side, the table in Figure 2 compares the time spent in the IPM solver on the CPU (using Pardiso ) and on the GPU (using the solver cuDSS). Overall, cuDSS is faster than Pardiso , leading to a 4x-8x speed-up in the total IPM solution time. We also note that the assembly of the condensed matrix $K_{\gamma}$ parallelizes well on the GPU, with a reduction in the assembly time from $\approx 8s$ on the CPU to $\approx 0.2s$ on the GPU.
Pardiso (CPU) cuDSS-$\mathrm{LDL^{T}}$ (CUDA)
$\gamma$ # it cond. (s) Cg (s) linsol (s) IPM (s) # it cond. (s) Cg (s) linsol (s) IPM (s)
$10^{4}$ 96 8.1 562.1 599.8 639.2 96 0.17 113.27 114.52 124.00
$10^{5}$ 96 8.5 212.0 249.9 290.0 96 0.17 53.37 54.62 64.39
$10^{6}$ 96 8.6 96.7 134.6 174.6 96 0.17 14.53 15.78 25.39
$10^{7}$ 96 8.6 52.9 91.0 130.5 96 0.17 7.95 9.20 18.41
$10^{8}$ 96 8.6 35.1 73.2 113.3 96 0.17 5.36 6.62 15.90
Figure 2: Above: Decomposition of IPM solution time across (a) condensation time (cond.), (b) Cg time, (c) total time spent in the linear solver (linsol.) and (d) total time spent in IPM solver (IPM). We compare the performance of cuDSS with those obtained by Pardiso running in parallel on the CPU using 8 threads. Below: Impact of $\gamma$ on the total number of Cg iterations and the norm of the relative residual at each IPM iteration. The peak observed in the norm of the relative residual corresponds to the primal-dual regularization performed inside the IPM algorithm, applied when the matrix $K_{\gamma}$ is not positive definite. The column for it displays the number of iterations, whereas the columns for cond., CG, linsol and IPM display the time spent in assembling the condensed KKT matrix, the time spent in the conjugate gradient, the time spent in the linear solver (including CG) and the total time spent in the IPM algorithm, respectively.
6.1.3 Tuning the equality relaxation strategy
We now analyze the numerical performance of LiftedKKT (§3.2). The method solves the KKT system (15) using a direct solver. The parameter $\tau$ used in the equality relaxation (14) is set equal to the IPM tolerance $\varepsilon_{tol}$ (in practice, it does not make sense to set a parameter $\tau$ below IPM tolerance as the inequality constraints are satisfied only up to a tolerance $\pm\varepsilon_{tol}$ in IPM).
We compare in Table 2 the performance obtained by LiftedKKT as we decrease the IPM tolerance $\varepsilon_{tol}$. We display both the runtimes on the CPU (using Pardiso with Cholesky) and on the GPU (using cuDSS-$\mathrm{LDL^{T}}$). The slacks associated with the relaxed equality constraints are converging to a value below $2\tau$, leading to highly ill-conditioned terms in the diagonal matrices $D_{s}$. As a consequence, the conditioning of the matrix $K_{\tau}$ in (15) can increase above $10^{18}$, leading to a nearly singular linear system. However, the factorization succeeds and the loss of accuracy caused by the ill-conditioning is tamed by the multiple Richardson iterations in the iterative refinement, reducing the relative accuracy in the residual down to an acceptable level. As a result, cuDSS can solve the problem to optimality in $\approx 20s$, a time comparable with HyKKT (see Figure 2). We observe that updating the parameter $\tau$ changes the total number of iterations in IPM: this is expected, as we solve a different problem (14) every time we update the regularization $\tau$.
Pardiso-$\mathrm{LDL^{T}}$ (CPU) cuDSS-$\mathrm{LDL^{T}}$ (CUDA)
$\varepsilon_{tol}$ #it time (s) #it time (s) accuracy
$10^{-4}$ 127 164.9 114 19.9 $1.2\times 10^{-2}$
$10^{-5}$ 129 168.1 113 30.4 $1.2\times 10^{-3}$
$10^{-6}$ 200 301.7 109 25.0 $1.2\times 10^{-4}$
$0^{-7} 11$0 159.2 104 20.1 $1.2\times 10^{-5}$
$0^{-8} 11$1 156.9 105 20.3 $1.2\times 10^{-6}$
Table 2: Performance of the equality-relaxation strategy as we decrease the IPM tolerance $\varepsilon_{tol}$. The table displays the wall time on the CPU (using Pardiso-$\mathrm{LDL^{T}}$) and on the GPU (using cuDSS-$\mathrm{LDL^{T}}$).
6.1.4 Breakdown of the time spent in one IPM iteration
We decompose the time spent in a single IPM iteration for LiftedKKT and HyKKT. As a reference running on the CPU, we show the time spent in the solver HSL MA27: The factorization time in HSL MA57 is 1.9s, compared to 1.2s for HSL MA27 (here, increasing the number of threads in the BLAS backend used by HSL MA57 does not make a difference). Hence, the block elimination algorithm implemented in HSL MA57 is not beneficial there.
When solving the KKT system, the time can be decomposed into: (1) assembling the KKT system, (2) factorizing the KKT system, and (3) computing the descent direction with triangular solves. As depicted in Figure 3, we observe that constructing the KKT system represents only a fraction of the computation time, compared to the factorization and the triangular solves. The condensed strategies LiftedKKT and HyKKT leads to faster factorization time, as the condensed KKT system can be factorized efficiently in parallel. HSL MA86 factorizes the condensed system almost 10 times faster than HSL MA27, and using cuDSS leads to an additional 4x speed-up. We focus exclusively on comparing the parallel CPU solvers (MA86 and Pardiso) for the Lifted/HyKKT formulations, as the factorization of the augmented KKT system ($K_{2}$) does not parallelize well on the CPU tasseff2019exploring . Once the KKT system is factorized, computing the descent direction with LiftedKKT is faster than with HyKKT (0.01s compared to 0.07s) as HyKKT has to run a Cg algorithm to completion to solve the Schur complement system (12) — hence requiring additional backsolves in the linear solver.
linear system linear solver build (s) factorize (s) backsolve (s) total (s)
($K_{2}$) HSL MA27 $2.85\times 10^{-2}$ $1.23\times 10^{0\phantom{-}}$ $3.48\times 10^{-1}$ $1.61\times 10^{0\phantom{-}}$
LiftedKKT (15) CHOLMOD $8.55\times 10^{-2}$ $5.84\times 10^{-1}$ $1.17\times 10^{-1}$ $7.86\times 10^{-1}$
LiftedKKT (15) pardiso $8.16\times 10^{-2}$ $2.23\times 10^{-1}$ $1.38\times 10^{-1}$ $4.43\times 10^{-1}$
LiftedKKT (15) MA86 $8.57\times 10^{-2}$ $36\times 10^{-1} 1.1$$18\times 10^{-1}$ $3.39\times 10^{-1}$
LiftedKKT (15) cuDSS $2.26\times 10^{-3}$ $4.91\times 10^{-2}$ $1.09\times 10^{-2}$ $6.23\times 10^{-2}$
HyKKT (11) CHOLMOD $8.01\times 10^{-2}$ $5.89\times 10^{-1}$ $7.22\times 10^{-1}$ $1.39\times 10^{0\phantom{-}}$
HyKKT (11) pardiso $9.06\times 10^{-2}$ $2.21\times 10^{-1}$ $9.19\times 10^{-1}$ $1.23\times 10^{0\phantom{-}}$
HyKKT (11) MA86 $8.60\times 10^{-2}$ $1.35\times 10^{-1}$ $7.50\times 10^{-1}$ $9.71\times 10^{-1}$
HyKKT (11) cuDSS $1.98\times 10^{-3}$ $3.38\times 10^{-2}$ $6.66\times 10^{-2}$ $1.02\times 10^{-1}$
Figure 3: Breakdown of the time spent in one IPM iteration for different linear solvers, when solving 78484epigrids. HSL MA27 factorizes the original augmented KKT system ($K_{2}$), and is used as a reference. The column for build shows the time spent in assembling the KKT matrix, whereas the columns for factorize and backsolve display the time spent in the numerical factorization and in the triangular solves, respectively. We set up Pardiso and HSL MA86 to run in parallel using 8 threads.
6.2 Benchmark on OPF instances
We run a benchmark on OPF instances from the PGLIB benchmark babaeinejadsarookolaee2019power . We compare the GPU solver configured with LiftedKKT and HyKKT against the CPU solver configured with the augmented KKT system and HSL MA27. For IPM solvers, we use the MadNLP solver shin2021graph with the ExaModels.jl modeling tool. We set the IPM tolerance to $10^{-6}$. Regarding HyKKT, we set $\gamma=10^{7}$, following the analysis in §6.1.2. The Dolan & Moré performance profile dolan2002benchmarking is displayed in Figure 4. The performance profile on the left compares LiftedKKT and HyKKT with MadNLP+Ma27. As a baseline, we include the performance obtained by Ipopt and Knitro, which also run with HSL MA27. Overall, the performance of HSL MA27 on the CPU is consistent with what was reported in babaeinejadsarookolaee2019power . The second performance profile on the right focuses on the performance obtained with LiftedKKT and HyKKT, using various sparse linear solvers. We observe that cuDSS leads to the fastest solution time compared to Panua Pardiso and HSL MA86 running in parallel on the CPU using 8 threads.
The detailed results are displayed in Table 3: The table displays the time spent on initialization, the time spent in the linear solver, and the total solving time. On the GPU, LiftedKKT+cuDSS is faster than HyKKT+cuDSS on small and medium instances; indeed, the algorithm does not have to run a Cg algorithm at each IPM iteration, limiting the number of triangular solves performed at each iteration. Both LiftedKKT+cuDSS and HyKKT+cuDSS are significantly faster than HSL MA27. HyKKT+cuDSS is slower when solving 8387_pegase, as in this particular instance, the parameter $\gamma$ is not set high enough to reduce the total number of Cg iterations, leading to a 4x slowdown compared to LiftedKKT+cuDSS. Nevertheless, the performance of HyKKT+cuDSS improves on the largest instances, with almost an 8x speed-up compared to the reference HSL MA27.
The benchmark presented in Table 3 was generated using NVIDIA A100 GPUs (80GB, current selling price: \$10k). We have also compared the performance with cheaper GPUs: a NVIDIA A1000 (a laptop-based GPU, 4GB memory) and a NVIDIA A30 (24GB memory, price: \$5k). As a comparison, the selling price of the AMD EPYC 7443 processor used for the benchmark on the CPU is \$1.2k. The results are displayed in Figure 5. We observe that the performance of the A30 and the A100 is relatively similar, and the cheaper A1000 GPU is already faster than HSL MA27 running on the CPU.
Among all nonlinear programs, the OPF problems have a very specific sparsity structure that leads to the formulation of a highly sparse KKT linear system ($K_{2}$). The super-sparse structure explains why HSL MA27 yields better results than HSL MA57, which can utilize parallelism at the BLAS level, as previously reported in tasseff2019exploring . For that reason, the numerical results obtained for the OPF instances may not be representative of the performance of HyKKT and LiftedKKT on generic nonlinear programs. In the next section, we analyze the performance of HyKKT and LiftedKKT on the CUTEst benchmark to provide a better overview of the methods’ limitations.
MadNLP+MA27 MadNLP+LiftedKKT+cuDSS MadNLP+HyKKT+cuDSS
(CPU) (GPU) (GPU)
it init AD lin total it init AD lin total it init AD lin total
1354_pegase 48 0.01 0.04 0.33 0.46 71 0.08 0.06 0.66 1.05 48 0.08 0.06 0.24 0.56
2000_goc 42 0.03 0.06 0.63 0.86 48 0.12 0.04 0.43 0.78 42 0.12 0.06 0.23 0.57
2312_goc 44 0.02 0.05 0.57 0.77 63 0.12 0.05 0.45 0.90 44 0.11 0.06 0.32 0.67
2742_goc 122 0.04 0.28 3.55 6.90 168 0.15 0.20 2.38 3.91 - - - - -
2869_pegase 55 0.05 0.11 1.00 1.42 57 0.15 0.05 0.33 0.80 55 0.14 0.08 0.40 0.88
3022_goc 55 0.03 0.10 0.92 1.26 53 0.13 0.06 0.30 0.74 55 0.13 0.07 0.37 0.81
3970_goc 48 0.06 0.14 1.89 2.64 48 0.22 0.05 0.49 1.00 48 0.20 0.09 0.40 0.92
4020_goc 59 0.06 0.18 3.56 4.46 64 0.22 0.05 0.77 1.60 59 0.21 0.12 0.67 1.33
4601_goc 71 0.06 0.22 3.08 4.11 67 0.21 0.06 0.60 1.22 71 0.21 0.14 0.67 1.40
4619_goc 49 0.07 0.18 3.07 3.95 52 0.26 0.05 0.70 1.34 49 0.25 0.10 0.54 1.16
4837_goc 59 0.06 0.20 2.44 3.39 56 0.22 0.05 0.44 1.10 59 0.22 0.11 0.57 1.21
4917_goc 63 0.06 0.19 1.85 2.75 69 0.20 0.06 0.79 1.51 63 0.20 0.10 0.54 1.16
5658_epigrids 51 0.08 0.21 2.69 3.71 48 0.27 0.05 0.55 1.23 51 0.26 0.11 0.56 1.22
7336_epigrids 50 0.11 0.26 3.44 4.64 54 0.36 0.06 1.12 2.08 50 0.37 0.11 0.58 1.36
8387_pegase 74 0.15 0.48 5.19 7.18 76 0.46 0.08 0.70 2.29 75 0.43 0.22 8.41 9.53
9241_pegase 73 0.38 0.53 5.71 8.07 107 0.50 0.12 1.18 3.42 70 0.49 0.14 1.68 2.76
9591_goc 67 0.41 0.49 10.11 12.37 75 0.49 0.09 1.99 3.74 67 0.48 0.15 1.23 2.27
10000_goc 82 0.16 0.50 5.65 7.69 64 0.39 0.07 1.12 2.21 82 0.35 0.19 1.06 2.09
10192_epigrids 54 0.42 0.46 7.18 9.42 68 0.57 0.08 1.54 3.14 54 0.51 0.14 1.08 2.09
10480_goc 71 0.40 0.60 11.01 13.68 66 0.60 0.08 1.57 3.51 71 0.56 0.14 1.61 2.74
13659_pegase 63 0.23 0.63 6.65 9.15 68 0.69 0.09 1.09 3.03 62 0.62 0.12 1.42 2.57
19402_goc 69 0.65 1.23 28.47 33.19 288 1.17 0.63 21.51 43.55 69 1.08 0.17 2.52 4.23
20758_epigrids 51 0.37 0.91 12.85 16.32 52 1.07 0.09 1.48 4.18 51 1.00 0.14 1.76 3.26
30000_goc 230 0.63 4.70 84.58 99.76 139 0.98 0.26 4.50 8.46 215 0.90 0.56 7.40 10.46
78484_epigrids 102 2.54 7.83 164.55 191.10 110 4.84 0.44 9.90 26.01 102 4.55 0.42 9.93 16.23
Table 3: OPF benchmark, solved with a tolerance tol=1e-6. The derivatives are evaluated using ExaModels.jl. The column for init displays the time spent initializing the IPM solver, including the initial symbolic factorization in the linear solver. The columns for AD, lin, and total display the time spent in the evaluation of the derivatives using AD, the time spent in the linear solver, and the total time spent in the IPM algorithm.
Refer to caption Refer to caption
Figure 4: Performance profile for the PGLIB OPF benchmark, solved with a tolerance tol=1e-6. Left: comparing LiftedKKT+cuDSS and HyKKT+cuDSS with HSL MA27. Right: comparing LiftedKKT and HyKKT with various parallel sparse linear solvers.
Figure 5: Comparing the performance obtained with various GPUs on three different OPF instances. We have used HyKKT to generate the results.
6.3 Benchmark on CUTEst instances
We have observed in the previous section that both LiftedKKT and HyKKT outperform state-of-the-art solvers running with HSL MA27 when running on the GPU. However, the OPF instances are specific nonlinear instances. In particular, the OPF instances exhibit extreme sparsity (with fewer than 10 non-zero coefficients per row, on average, in the augmented KKT system). Therefore, they cannot be considered representative nonlinear programs. For that reason, we complement our analysis by looking at the performance of LiftedKKT and HyKKT on large-scale CUTEst instances. Since the parallel solution approaches are intended for large-scale problems, we select the instances with more than 1,000 variables and at least 1 constraint.
As for the PGLIB benchmark, we set the IPM tolerance to $10^{-6}$ and uses $\gamma=10^{7}$ in HyKKT. We use the package CUTEst.jl to decode the SIF instances on the CPU. To use CUTEst on the GPU with HyKKT+cuDSS and Lifted+cuDSS, we have implemented a wrapper moving the sensitivities (gradient, constraints, Jacobian and Hessian) on the GPU using a copy. As a result, the evaluation of the derivatives is not as fast as if they were evaluated on the GPU, and the GPU solvers are slightly penalized by the data exchange between the CPU and the GPU.
A performance profile comparing LiftedKKT+cuDSS and HyKKT+cuDSS is provided in Figure 6. As a baseline, we include the performance obtained with Ipopt, MadNLP, and Knitro, all running with HSL MA57 (in contrast to OPF problems, HSL MA57 has better performance than HSL MA27 on CUTEst). In the performance profile (a), we observe that HyKKT and LiftedKKT are less competitive than a standard method running on the CPU, despite demonstrating robustness comparable to Knitro for LiftedKKT. However, if we select only the largest instances where Ipopt solve time is greater than 1 second (performance profile (b)), HyKKT and LiftedKKT become more competitive and are overall as fast as Knitro (the best-performing solver on the largest CUTEst instances) on 50% of the instances.
Overall, HyKKT is less robust than LiftedKKT on CUTEst: in some cases, the regularization parameter $\gamma=10^{7}$ is not set high enough to ensure a positive definite matrix $K_{\gamma}$. For instance, this is the case on ROCKET: this optimal control is known to be degenerate, in the sense that the reduced Hessian $Z^{\top}KZ$ is close to being singular (this is induced by the infamous singular arc in Goddard’s problem). We have observed that $\gamma$ must be increased to $10^{12}$ to obtain a positive definite matrix $K_{\gamma}$.
Figure 6 shows the two condensed space methods are less robust than the reference methods running on the CPU. To analyze further their performance, we select in Table 4 a dozen of representative instances where the condensed space methods are either significantly slower (unfavorable cases) or significantly faster than the references (favorable cases). Overall, the unfavorable cases correspond to one of the following situations. (i) The linear solver cuDSS takes significant time computing the symbolic factorization, as it is the case on BLOCKQP2 or ROSEPETAL2. (this explains the bump observed at the upper right of the performance profiles in Figure 6). (ii) The Jacobian has a dense column (ROCKET, BRAINPC0, BLOCKQP2, ROSEPETAL2) or almost dense columns (A5NSDSIL, A0NNDNIL, A0NSDSIL, A0NNSNSL, MPC9, CMPC10, CMPC12, DALE, CBS), leading to additional fill-in when factorizing the condensed matrices $K_{\gamma}$ and $K_{\tau}$. (iii) the equality relaxation strategy used in LiftedKKT results in a poor convergence in IPM (LUKVLE8, READING7). However, we observe on the favorable cases that if the problem does not suffer from one of the previous pitfalls, HybridKKT and LiftedKKT become competitive w.r.t. the references, with speed-up similar to what we observed before on the OPF instances (Table 3). We observe that the condensed methods are efficient to solve problems with a large number of non-zero coefficients in the Jacobian and in the Hessian (BDRY2, CLEUVEN*) or problems with a dynamical structure (MARINE, ROBOTARM). Interestingly, we observe that the condensed-space methods can efficiently accommodate dense condensed matrices $K_{\tau}$ and $K_{\gamma}$ provided they remain small enough (EIGENAU, EIGENCCO, EIGENB2, MSQRTB, where the size of the dense condensed matrix is at most $\approx 2,500\times 2,500$, compared to more $10,000\times 10,000$ in the unfavorable cases).
Refer to caption Refer to caption
(a) (b)
Figure 6: Performance profiles for the CUTEst benchmark, selecting all instances with more than 1,000 variables and at least 1 constraint. The profile (a) displays the entire benchmark, whereas the profile (b) selects only the largest instances where Ipopt takes more than 1s to run.
HSL MA57 LiftedKKT+cuDSS HyKKT+cuDSS
n m nnzj nnzh it init AD lin total it init AD lin total it init AD lin total
GPU-accelerated solvers are significantly slower
A5NSDSIL 60012 20004 115012 100020 265 0.03 1.22 6.97 11.33 119 26.16 1.35 129.95 164.00 - - - - -
A0NNDNIL 60012 20004 115003 100020 97 0.02 0.43 2.22 3.15 470 17.39 4.10 365.43 395.94 97 17.18 0.93 73.82 93.57
A0NSDSIL 60012 20004 115005 100020 100 0.02 0.43 2.23 3.15 212 18.41 1.88 168.46 192.80 100 18.50 1.20 79.08 100.67
A0NNSNSL 60012 20004 111099 100020 31 0.04 0.14 1.36 1.62 36 19.00 0.34 38.76 58.78 38 19.24 0.48 80.65 101.15
CMPC10 1530 2351 16538 357 70 0.00 0.02 0.24 0.27 82 0.10 0.29 1.93 4.90 70 0.04 0.06 0.78 1.07
CMPC12 1530 2351 16538 357 78 0.00 0.02 0.27 0.31 99 0.21 0.79 4.81 13.11 77 0.04 0.07 0.82 1.14
MPC9 1530 2351 16538 357 70 0.00 0.02 0.25 0.29 79 0.21 0.63 3.25 9.78 69 0.04 0.07 0.75 1.05
LUKVLE8 10000 9998 29994 30000 12 0.01 0.05 0.04 0.10 144 0.12 1.36 2.48 25.83 12 0.04 0.05 13.45 13.57
READING7 1002 500 126251 1503 125 0.01 0.14 1.28 1.49 330 0.65 1.45 4.75 18.18 120 0.55 0.35 1.25 3.09
BRAINPC0 6907 6900 41399 17335 120 0.01 0.27 0.78 1.11 43 0.22 0.47 2.83 7.65 - - - - -
ROCKET 2407 2002 10405 8019 58 0.00 0.05 0.10 0.16 38 0.21 0.35 1.71 5.11 - - - - -
HYDROELL 1009 1008 2016 2017 235 0.00 0.03 0.08 0.15 235 0.05 0.41 0.70 4.65 235 0.02 0.16 0.19 0.95
DALE 16514 405 33028 16514 14 0.01 0.01 0.07 0.11 16 2.12 0.05 30.59 33.03 14 1.66 0.03 16.60 18.36
CBS 11163 244 22326 11163 20 0.00 0.01 0.07 0.10 22 0.75 0.03 7.60 8.49 35 0.74 0.07 8.92 9.91
BLOCKQP2 10010 5001 70010 15010 19 0.01 0.02 0.09 0.13 37 36.27 0.08 7.37 44.62 19 36.30 0.05 2.75 39.50
ROSEPETAL2 10001 20001 50001 10000 29 0.05 0.04 0.15 0.29 47 36.11 0.17 7.41 44.86 29 36.15 0.17 3.24 40.16
GPU-accelerated solvers are significantly faster
STCQP1 8193 4095 28665 57333 13 0.01 0.06 13.87 13.96 15 0.06 0.08 0.06 0.26 161 0.07 1.10 1.18 2.94
CVXQP3 10000 7500 22497 39984 17 0.12 0.03 36.33 36.50 31 0.13 0.07 0.41 0.69 17 0.13 0.07 118.33 118.60
CATENA 3003 1000 6000 6003 32 0.00 0.02 12.72 12.74 36 0.06 0.10 0.33 1.13 - - - - -
CLEUVEN3 1200 2973 199600 71583 134 0.08 0.80 6.81 7.87 104 0.53 0.80 0.29 2.16 99 0.54 0.84 1.07 3.00
NCVXQP7 10000 7500 22497 39984 121 0.12 0.21 133.73 134.46 302 0.13 1.12 7.91 10.55 129 0.13 0.48 178.30 179.46
BDRY2 251001 250498 1247999 63500 45 2.53 1.06 158.10 163.68 157 1.46 4.99 28.96 36.30 50 1.75 1.57 10.99 14.77
SCW2 2003 1001 2002 2007006 251 0.41 25.96 66.30 95.12 97 0.75 10.64 1.55 13.48 259 0.74 28.03 4.71 35.14
CLEUVEN6 1200 3091 216608 71583 206 0.08 1.26 9.45 11.05 88 0.55 0.74 0.27 2.04 - - - - -
AUG3D 27543 8000 50286 22743 2 0.16 0.01 1.45 1.61 3 0.22 0.01 0.04 0.29 1 0.22 0.01 0.03 0.26
ROBOTARM 4412 3202 15605 8422 451 0.00 0.28 1.38 1.81 36 0.06 0.04 0.18 0.46 - - - - -
EIGENB2 2550 1275 125000 66300 132 0.08 0.82 211.36 212.42 110 1.23 1.01 1.77 4.68 128 1.23 1.21 2.14 5.25
CBRATU3D 3456 2000 16000 3000 3 0.03 0.00 0.29 0.32 5 0.06 0.01 0.03 0.13 3 0.07 0.01 0.03 0.12
EIGENC2 2652 1326 132651 70278 16 0.08 0.10 26.37 26.57 17 1.31 0.16 0.58 2.15 15 1.32 0.16 0.56 2.12
STNQP1 8193 4095 28665 57333 17 0.01 0.07 19.59 19.69 21 0.06 0.12 0.13 0.38 60 0.06 0.40 0.48 1.24
EIGENCCO 2652 1326 132651 3517878 42 0.58 8.85 51.38 61.37 35 1.46 7.56 0.59 9.84 42 1.77 9.09 0.83 11.93
EIGENAU 2550 2550 313750 66300 2 0.05 0.03 10.95 11.05 3 1.49 0.05 0.02 1.60 1 1.50 0.03 0.01 1.57
LUKVLE12 9997 7497 19992 22492 86 0.00 0.18 208.54 208.77 30 0.05 0.09 0.09 0.41 24 0.04 0.07 0.08 0.30
MSQRTB 1024 1024 64512 33264 5 0.01 0.01 1.62 1.65 6 0.23 0.02 0.03 0.30 5 0.23 0.02 0.02 0.29
DRCAVTY1 4489 3969 51597 86879 8 0.01 0.07 1.19 1.28 8 0.08 0.09 0.05 0.25 7 0.08 0.09 0.05 0.24
CLEUVEN5 1200 2973 199600 71583 134 0.07 0.81 6.80 7.85 104 0.53 0.82 0.29 2.18 99 0.53 0.77 1.11 3.04
MARINE 11215 11192 46768 9423 85 0.01 0.11 8.60 8.75 25 0.09 0.05 0.14 0.39 31 0.10 0.07 0.25 0.54
Table 4: CUTEst benchmark, solved with a tolerance tol=1e-6. The derivatives are evaluated with CUTEst. The column init displays the time spent initializing the IPM solver, including the initial symbolic factorization in the linear solver. The columns AD, lin and total display the time spent in the evaluation of the derivatives using AD, the time spent in the linear solver and the total time spent in the IPM algorithm.
7 Conclusion
This paper investigates the theoretical and numerical properties of two condensed-space interior-point methods for solving large-scale nonlinear programming problems on GPUs. Despite the emergence of ill-conditioned matrices, our theoretical analysis shows that the resulting loss of accuracy is benign in floating-point arithmetic, owing to the structural properties of the interior-point method. Numerically, both methods prove competitive for solving large-scale nonlinear problems. The fully GPU-resident software stack we developed, which includes MadNLP.jl, cuDSS, and ExaModels.jl, demonstrates its effectiveness on large-scale optimal power flow (OPF) problems from the pglib-opf library. Notably, this GPU-based implementation achieves up to a 10× speedup on large-scale OPF problems. While performance is more variable across the CUTEst benchmarks, our solver consistently remains competitive with its CPU-based counterpart. When focused on the instances with non-trivial size, the GPU implementation achieves considerable speedups, at the price of reduced robustness in the linear solver. These results highlight both the potential (high performance and scalability) and the limitations (robustness and convergence) of GPU-based interior-point methods for nonlinear programming. Looking ahead, we plan to improve the robustness of the two condensed KKT methods, with a particular emphasis on achieving stable convergence at tighter tolerances (below $10^{-8}$). Moreover, the sparse Cholesky solver can be further customized to align with the specific structure and needs of interior-point algorithms wright1999modified .
Fundings
This research used resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357.